### v2

# MMAR: Towards Lossless Multi-Modal Auto-Regressive Probabilistic Modeling

Abstract
Recent advancements in multi-modal large language models have propelled the development of joint probabilistic models capable of both image understanding and generation. However, we have identified that recent methods inevitably suffer from loss of image information during understanding task, due to either image discretization or diffusion denoising steps. To address this issue, we propose a novel Multi-Modal Auto-Regressive (MMAR) probabilistic modeling framework. Unlike discretization line of method, MMAR takes in continuous-valued image tokens to avoid information loss. Differing from diffusion-based approaches, we disentangle the diffusion process from auto-regressive backbone model by employing a light-weight diffusion head on top each auto-regressed image patch embedding. In this way, when the model transits from image generation to understanding through text generation, the backbone model’s hidden representation of the image is not limited to the last denoising step. To successfully train our method, we also propose a theoretically proven technique that addresses the numerical stability issue and a training strategy that balances the generation and understanding task goals. Through extensive evaluations on 18 image understanding benchmarks, MMAR demonstrates much more superior performance than other joint multi-modal models, matching the method that employs pretrained CLIP vision encoder, meanwhile being able to generate high quality images at the same time. We also showed that our method is scalable with larger data and model size. Code will be available at https://github.com/ydcUstc/MMAR.
1 Introduction
Over the past few years, extensive research in the field of multimodal intelligence has catalyzed the accelerated advancement of foundational models for both image understanding and image generation. Within the realm of image understanding, multimodal large language models (MLLM), exemplified by LLaVA (Liu et al., 2023a), have exhibited human-like capabilities in open-domain image comprehension. In the domain of image generation, techniques rooted in generative probabilistic models, such as Denoising Diffusion Probabilistic Models (DDPM) (Ho et al., 2020) and auto-regressive (AR) models (Chen et al., 2020), have also garnered significant success. Essentially, these two lines of research correspond to modeling the conditional probability, i.e., $P(T|I)$ and $P(I|T)$, where $T$ and $I$ corresponds to text and image, respectively. It’s evident that both types of conditional probabilistic models are subsets of a joint probabilistic model, $P(T,I)$. This brings us to an intriguing question: Could a joint probabilistic model serve as a natural unifying framework for both image understanding and image generation tasks?
Given that the most advanced image understanding (Chen et al., 2024) and generation (Esser et al., 2024) models rely on the language priors $p(T)$ of pre-trained large language models (LLMs), the most straight-forward approach for joint image-text probabilistic modeling is to convert images into discrete tokens similar to text. This way, images are treated as a form of language and integrated into the auto-regressive modeling of text, as seen in models like MARS (He et al., 2024), LlamaGen (Sun et al., 2024a), and Chameleon (Team, 2024). This method leverages the powerful text modeling capabilities of various open-source pre-trained LLMs, along with their highly optimized training and inference frameworks. However, discretization on the continuous image space introduces an information bottleneck, inevitably leading to a loss of image details and reducing the model’s information capacity. This limitation is quantitatively evident in the image understanding performance: existing methods based on image token discretization fall short when compared to the LLaVA model, which utilizes continuous CLIP representations (Xie et al., 2024).While there are theoretically feasible methods to mitigate this information bottleneck, such as significantly increasing the number of tokens or the size of the VQ codebook, these approaches would substantially increase the training difficulty for both LLMs and VQ-VAE models.
Recently, several attempts have been made to address the continuous nature of images by combining image diffusion transformers and text auto-regression within a unified transformer architecture (Zhou et al., 2024; Zhao et al., 2024). Although this approach, exemplified by Transfusion (Zhou et al., 2024) and MonoFormer, demonstrated superior performance compared to discrete token auto-regressive methods for images, it is crucial to acknowledge that the inherent nature of image diffusion prevents this approach from leveraging complete image modeling capability when it comes to image understanding tasks, as shown in fig. 1-(B). The root cause is that image diffusion will encode different parts of image information into different noise levels (Ho et al., 2020). For example, at low noise level, the model tends to solve an image enhancement task, while at high noise level, the model is only suitable for extracting rough image morphology and layout. Intuitively, image-to-text tasks need to take clean images as input to maintain complete visual information. However, due to the nature of diffusion modeling, this will only invoke the hidden representations that are related to the image enhancement task according to its low noise level, resulting in insignificant help for image-to-text tasks. Adding some noise to the input image would help invoke more semantic representations, but at a cost of image information loss. To ensure a full utilization of the image modeling capacity learned by image diffusion, it is necessary to input the images of all noise levels, but with tremendous computational overhead. This dilemma hinders transfusion-like models perform well for visual understanding tasks.
Figure 1: Comparison of different multi-modal probabilistic modeling paradigms. (A): Most AR-based methods adopt discrete image tokens produced by VQ-VAEs. The VQ operation bottlenecks the representation’s information capacity, resulting in information loss. (B): Concurrent transfusion-like methods combining image DiT with text AR. $X^{(i)}$ denotes the full image at noise level $i$. For DiT part, the image information is modeled within the learned representation across various noise level, represented by $h^{(0)},...,h^{(t)}$, where the superscript indicates noise level. However, only $h^{(0)}$ is utilized in image to text tasks. (C): Our MMAR paradigm is VQ-free, thus reducing the information loss related to VQ operation. It auto-regressively models the continuous image tokens $x_{1},...,x_{n}$. The auto-regressive nature guarantees that all the image hidden representations, denoted by $h_{i},i\in[1,n]$, are simultaneously optimized for both $P(I)$ and $P(T|I)$, ensuring complete usage of image modeling capability for visual understanding tasks.
In summary, how to take in and utilize the complete information of the continuous image modality is the major pain point of joint image-text probabilistic modeling. This is the challenge that our work is trying to address. The method proposed in this paper, MMAR, belongs to the image-text joint AR methodology, thus naturally avoids the challenge of fusing diffusion and AR, as mentioned previously. Different from other image-text joint AR methods, our method is built upon continuous image representation that can faithfully reconstruct the original image, rather than the discretized lossy image representation. To model the continuous probability distribution within the auto-regressive paradigm, we refer to MAR (Li et al., 2024b), introducing a simple mlp diffusion sampler to sample continuous image tokens given the transformer backbone output as its conditional information. This technique can also be regarded as disentangling the diffusion process from the auto-regressive backbone in the transfusion-like methods (Zhou et al., 2024; Zhao et al., 2024), which in turn focuses on learning complete visual representation that can guide the entire diffusion sampling process of image tokens. Different from MAR (Li et al., 2024b), we leverage LLMs to learn much more diverse distribution of the large-scale image-text data. In practice, we find that low-precision training strategies for contemporary LLMs introduce non-negligible numerical error term in the diffusion sampler, which substantially impedes the convergence. We carefully analyze the numerical error term and find that the only way to minimize it is to properly parameterize the diffusion model. Therefore, we theoretically derive an optimal diffusion parameterization to solve this issue. Additionally, we find that visual understanding and generation prefer different experimental settings. To balance these two tasks in a unified model, we propose a two-stage training method, which firstly improves visual understanding ability with large, mid quality data, then boost visual generation ability with high quality data. With these innovations, the lossless multi-modal auto-regressive probabilistic modeling is finally achieved theoretically and practically.
Our contributions are the following three folds:
  • We identified why previous joint probabilistic models for both image generation and understanding suffer image information loss with a series of experiments.
  • We proposed a novel modeling framework that solves the issue by disentangling diffusion process from auto-regressive backbone, and achieved significant performance improvement for image understanding task.
  • We proposed two training techniques crucial for training the model: one solves numerical error issue under low-precision training setting with theoretical guarantee, the other balances image generation and understanding tasks goals.
2 Related Works
2.1 Multi-modal Large Language Models
Since LLMs demonstrated open-domain conversational capabilities, more and more research has be focused on how to introduce visual information into large language models to achieve open-domain visual understanding. Pioneering works such as BLIP-2 (Li et al., 2023b), MiniGPT4 (Zhu et al., 2024) and LLaVA (Liu et al., 2023a) use trainable connector modules such as qformer or mlp to align image representations to the input space of LLM, making open-domain visual question answering possible. Recently, thanks to innovations in network architecture (Dong et al., 2024b; Tong et al., 2024), training strategy (Chen et al., 2024) and the support for dynamic resolution input (Hu et al., 2024; Li et al., 2024a), the visual understanding performance of large multi-modal models have been greatly improved. These works focus on the alignment of image representations to text representations, only achieving $p(T|I)$ without incorporating the image’s own distribution $p(I)$ into the modeling capabilities of the model. Different from these works, our work adds the modeling of $p(I)$ on the basis of these MLLMs to achieve joint image-text probabilistic modeling.
2.2 Auto-Regressive Image Generative Models
Text-to-image generative models aim at modeling the conditional probability distribution $p(I|c)$, enabling probabilistic sampling of images conditioned on textual or categorical inputs. Auto-regressive methods represent a dominant paradigm in this domain, typically requiring discrete representations for both input and output. For images, this necessitates encoding them into discrete codes using a VQVAE (Esser et al., 2021; Kondratyuk et al., 2024). While recent works demonstrate that auto-regressive methods based on discrete image tokens can generate high-quality images (Sun et al., 2024a), the discretization of image representation acts as an information bottleneck, limiting the modeling accuracy of the image distribution. Recent efforts have shown that auto-regressive probabilistic modeling can be achieved without relying on discrete representations (Tschannen et al., 2023; Li et al., 2024b). For instance, MAR (Li et al., 2024b) replaces traditional logits with diffusion heads, enabling probabilistic sampling of continuous representations within an auto-regressive framework. This paper introduces continuous representation auto-regressive probability modeling to MLLMs, mitigating information loss caused by quantization and achieving theoretically lossless joint image-text probability modeling. In addition, we addressed the difficulty when training with large model and large data which is not presented in the previous works.
2.3 Unified Visual Understanding and Generation Models
Recently, a series of studies have focused on leveraging a single model to simultaneously address tasks of visual generation and understanding. Early works in this area adopted a modular approach, bridging pre-trained models for visual understanding and visual generation through intermediate representations to achieve combined image understanding and generation. Notable examples include EMU (Sun et al., 2024c, b) and SEED-X (Ge et al., 2024). These works, however, are not considered probabilistic modeling because they aim at modeling the mean value of representations like CLIP or other intermediate representations rather than modeling the true image distribution. This limitation leads to the inadequate image space exploration, and thus hinders the attainment of high-quality generative and understanding performance.
Another line of research adheres to the paradigm of probabilistic modeling (Team, 2024; He et al., 2024; Xie et al., 2024; Wu et al., 2024; Zhou et al., 2024; Zhao et al., 2024). These approaches can be categorized into three types based on whether the image representations are discrete and the modeling method of the image part : (i) Discrete auto-regressive Methods: Examples include Chameleon (Team, 2024), MARS (He et al., 2024), and VILA-U (Wu et al., 2024). These methods discretize image representations and then model images and text tokens using a unified auto-regressive transformer. (ii) Discrete Diffusion Methods: An example is Show-o (Xie et al., 2024). These methods discretize image tokens and model them with text tokens using a unified transformer through a discrete diffusion approach. (iii) Continuous Diffusion Methods: Examples include Transfusion (Xie et al., 2024) and MonoFormer (Zhao et al., 2024). These methods do not discretize image representations but directly employ continuous diffusion transformers to model image tokens along with text tokens using a unified transformer. Our approach differs from the aforementioned three types. It belongs to the continuous auto-regressive method category, which does not require discretizing image representations. Instead, it predicts the continuous distribution of image tokens using an auto-regressive approach and models them alongside text tokens within a unified transformer, as shown in fig.1.
3 Method
3.1 Auto-Regressive Modeling with Continuous and Discrete Representations
Auto-regressive modeling is a commonly used probabilistic modeling method for sequence data. It can be formulated by “predicting the next token” as follows:
$\log p_{\theta}(\mathbf{x})=\sum_{i=1}^{n}\log p_{\theta}(x_{i}|x_{<i}),$ (1)
where $\theta$ and $\mathbf{x}$ represent model parameters and the sequence data, respectively. By maximizing the log likelihood of the data, $\mathbb{E}_{\mathbf{x}\sim\mathcal{D}}\log p_{\theta}(\mathbf{x})$, the model can be optimized to sample from the data distribution $\mathcal{D}$, achieving probabilistic modeling.
In the realm of natural language processing (NLP), the sequence $\mathbf{x}$ is solely made of discrete text tokens. As a result, most modern large language models (LLMs) parameterize $p_{\theta}(x_{i}|x_{<i})$ into a categorical distribution, which can be explicitly represented by the softmax activation on a set of logits predicted by a decoder-only transformer (Radford et al., 2019; Brown et al., 2020) $f_{\theta}(\cdot)$ followed by a linear LM head $H_{\theta}(\cdot)$:
$p_{\theta}(x_{i}|x_{<i})=\mathrm{softmax}(H_{\theta}(f_{\theta}(x_{<i}))).$ (2)
In addition to text, our work also aims at modeling the probability of images, which are represented by continuous rather than discrete image tokens. Therefore, a protocol for parameterizing $p_{\theta}(x_{i}|x_{<i})$ of the continuous image tokens is required. Inspired by MAR (Li et al., 2024b), we train a diffusion model to achieve this. The diffusion model takes vector $z_{i}=f_{\theta}(x_{<i})$ as the conditional input, and samples $x_{i}$ by gradually denoising from a randomly sampled gaussian noise. To optimize the diffusion model for continuous image token sampling, a diffusion loss can be applied, which acts as the upper-bound of the negative log likelihood. A typical diffusion loss can be written as follows, which is seen in MAR (Li et al., 2024b):
$L(x_{i})=\mathbb{E}_{x_{i},\epsilon,t}[w_{t}\cdot||\mathbf{\epsilon}-\mathbf{\epsilon}_{\theta}(\sqrt{\bar{\alpha}_{t}}x_{i}+\sqrt{1-\bar{\alpha}_{t}}\mathbf{\epsilon},t,z_{i})||^{2}]\geq-\log p_{\theta}(x_{i}|x_{<i})+C,$ (3)
where $w_{t}$ is the loss weight that balances the loss for different timesteps, and $\bar{\alpha}_{t}$ indicates the noise schedule of the forward diffusion process. In this way, minimizing the diffusion loss is equivalent to maximizing the log likelihood of image data.
The overall loss for joint image-text probabilistic modeling can be written as follows:
$L=\sum_{i\in I_{img}}L(x_{i})-\sum_{i\in I_{txt}}\log p_{\theta}(x_{i}|x_{<i}),$ (4)
where $I_{img}$ and $I_{txt}$ indicate the indices of image tokens and text tokens, respectively.
3.2 Optimal Diffusion Model Parameterization for Low-Precision Training
In the era of large language models, training with low precision data type, such as bfloat16, has become increasingly popular. However, the training and inference process of a diffusion model is relatively sensitive to numerical precision. Moreover, in an auto-regressive framework, the image tokens are sampled sequentially, requiring even more precise sampling for each image token to reduce the overall error accumulation. Therefore, handling the numerical error in the diffusion process modeling should be emphasized when integrating diffusion loss into LLMs.
From the example below, we can clearly illustrate the effect of the numerical error in diffusion models. In eq.3, the diffusion model is parameterized as $\mathbf{\epsilon}_{\theta}(\sqrt{\bar{\alpha}_{t}}x_{i}+\sqrt{1-\bar{\alpha}_{t}}\mathbf{\epsilon},t,z_{i})$, known as “$\epsilon$-prediction”, predicting the noise $\epsilon$ that is added to the data. Floating-point representation causes numerical error’s amplitude proportional to a number’s magnitude. This can be modeled by scaling the prediction by a factor of $(1+\delta)$, where $\delta$ is the relative error. In this way, we can write DDIM (Song et al., 2021) sampling with numerical error as follows:
$\tilde{x}^{(t-1)}=$ $\sqrt{\bar{\alpha}_{t-1}}\left(\frac{x^{(t)}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(x^{(t)},t,z_{i})(1+\delta)}{\sqrt{\bar{\alpha}_{t}}}\right)+\sqrt{1-\bar{\alpha}_{t-1}}\epsilon_{\theta}(x^{(t)},t,z_{i})(1+\delta).$ (5)
Further separating the numerical error term from the ideal DDIM sampling process, we get:
$\tilde{x}^{(t-1)}=x^{(t-1)}+(\sqrt{1-\bar{\alpha}_{t-1}}-\frac{\sqrt{\bar{\alpha}_{t-1}}}{\sqrt{\bar{\alpha}_{t}}}\sqrt{1-\bar{\alpha}_{t}})\epsilon_{\theta}(x^{(t)},t,z_{i})\delta,$ (6)
where the first term $x^{(t-1)}$ is the ideal DDIM sampling term, and the second term is the numerical error term. Assuming $\epsilon_{\theta}(x^{(t)},t,z_{i})$ follows a standard normal distribution, the standard deviation of numerical error term can be calculated as $|\sqrt{1-\bar{\alpha}_{t-1}}-\frac{\sqrt{\bar{\alpha}_{t-1}}}{\sqrt{\bar{\alpha}_{t}}}\sqrt{1-\bar{\alpha}_{t}}|\delta$. When the signal-to-noise ratio (SNR) is high, i.e. $\bar{\alpha}_{t},\bar{\alpha}_{t-1}\to 1$, the numerical error has almost zero amplitude. However, when SNR is extremely low, i.e. $\bar{\alpha}_{t}=0$, and $\bar{\alpha}_{t-1}>0$, this term will explode to infinity, causing extreme numerical error.
Our goal is to minimize the numerical error term. To achieve this, we analyze all the factors that can determine the numerical error in Appendix A.2, and find that the only solution is to parameterize the diffusion model properly. By solving the numerical error minimization problem, we conclude that the v-prediction parameterization (Salimans & Ho, 2022) is the desired optimal parameterization method. Note that v-prediction is initially proposed for the efficient distillation of diffusion models, rather than reducing the numerical error of diffusion models. To the best of our knowledge, our work is the first to derive v-prediction parameterization from the first principle of minimizing the numerical error in diffusion models. For more details, see Appendix A.2. Under v-prediction parameterization, the model predicts a linear combination of data $x_{i}$ and noise $\epsilon$:
$v^{(t)}_{i}=\sqrt{\bar{\alpha}_{t}}\mathbf{\epsilon}-\sqrt{1-\bar{\alpha}_{t}}x_{i}$ (7)
We therefore re-write eq.3 into the v-prediction form, and set the loss weight $w_{t}$ to 1 for simplicity:
$L(x_{i})=\mathbb{E}_{x_{i},\epsilon,t}[||v_{i}^{(t)}-v_{\theta}(\sqrt{\bar{\alpha}_{t}}x_{i}+\sqrt{1-\bar{\alpha}_{t}}\mathbf{\epsilon},t,z_{i})||^{2}].$ (8)
3.3 Model Architecture and Training Strategy
Figure 2: The overview of MMAR with two stage image expert training strategy.
3.3.1 Pipeline
To jointly model the probabilities of images and text at the token level, we employ an image tokenizer to extract image tokens and randomly mask some of them for image generation training, as depicted in fig. 2. Given the substantial distribution gap between images and text, directly scaling image tokens to match the LLM’s dimensions would pose a challenge to joint image-text probability modeling. To address this, we introduce a Latent Encoding process utilizing an EmbeddingViT module for the unmasked image tokens, which simplifies the modeling task. To facilitate the auto-regressive training of image tokens, we concatenate the known image tokens with the masked tokens and append the position embedding corresponding to their original positions. These tokens, after being processed by the VisProjector, are concatenated with the text tokens to generate the final image-text sequence. Then, an LLM is used to process the image-text sequence, whose output $z$ integrates information from both the preceding text and the unmasked image tokens. For image output, $z$ acts as a conditional input for the Diffusion MLP, which predicts the $v$ values of the noisy masked image tokens. The diffusion loss, as depicted in eq.8, is applied, enabling us to model the probability of the continuous image tokens. For text output, a linear lm-head processes $z$ to predict the next token’s logits, and the corresponding cross entropy loss is applied, allowing us to model the probability of the text tokens.
During image generation, to ensure consistency with the training scenario, we initialize all image tokens as mask tokens and randomly generate a set of image position sequences. Following a left-to-right order, we extract N condition tokens $z$ at each step and input them into the Diffusion MLP to generate the corresponding image tokens. Next, we assign these image tokens back to their respective positions in the original image token sequence, iterating until the entire image token set is obtained. Ultimately, we decode the image tokens into images using the image tokenizer, yielding the generated images.
3.3.2 Model Architecture
Tokenizer
MMAR employs the publicly available KL-16 image tokenizer from LDM (Rombach et al., 2022). This tokenizer processes an 256 $\times$ 256 image into 256 image tokens, each with 16 channels, and it is kept frozen during training.
EmbeddingViT
To achieve a more profound encoding of image tokens, we introduce the EmbeddingViT module, a ViT(Dosovitskiy et al., 2021) encoder with 16 layers and 1024 hidden state channels, processing image tokens into visual embeddings with stronger context awareness. Moreover, as the input consists of randomly shuffled image tokens, we integrate a learnable position embedding for each image token in EmbeddingViT, corresponding to its original position.
LLM
We initialize our LLM using parameters from the open-source QWen2 series models (Yang et al., 2024). To preserve text modeling capabilities, we keep the LLM parameters fixed during training and add PLoRA (Dong et al., 2024b) as the image expert, where only the image tokens in the input pass through the introduced LoRA (Hu et al., 2022) adapters. The PLoRA is applied for each linear layer within the original LLM. Considering the time cost, our ablation study employs QWen2-0.5B-Instruct, with $r=512$ for PLoRA. Furthermore, we use QWen2-7B-Instruct to explore our method’s scale-up capability with $r=1280$ for PLoRA. The corresponding MMAR models are denoted by MMAR-0.5B and MMAR-7B, respectively.
In the LLM, we adopt a bidirectional attention mechanism to enhance interaction among image tokens, rendering all image tokens mutually visible, as depicted in fig. 2. Meanwhile, the text portion retains causal attention. Additionally, to prevent interference with the auto-regressive training of image tokens in random order, we assign the same ROPE position id to all image tokens within all LLM layers. This strategy not only ensures the random order regression of image tokens but also mitigates the issue of text concentrating on closer image tokens in long text scenarios.
Diffusion MLP
Inspired by MAR, we employ a simple MLP to predict $v^{(t)}$. This MLP consists of a series of residual blocks, each comprising AdaLN (Peebles & Xie, 2023), a linear layer, SiLU, and another linear layer. The condition vector $z$ is added to the diffusion time embedding and incorporated through AdaLN. MMAR-0.5B utilizes a Diffusion MLP with 8 residual blocks and 1024 channels, while MMAR-7B employs a larger variant with 12 residual blocks and 2048 channels.
3.3.3 Training Strategy
Multi-Task Training
To accomplish joint image-text modeling, we simultaneously conducted text-to-image, image-to-text, and unconditional image generation tasks during training. In the image-to-text task, no mask tokens are assigned to the image part, allowing us to model $P(T|I)$ using the complete image tokens. Furthermore, the allocation ratio of text-to-image and unconditional image generation tasks is set to 9:1, facilitating efficient use of the cfg technique at the inference stage. To maintain a balance between tasks, we intuitively set the sample allocation ratio of image generation tasks and image understanding tasks to 1:1.
Two Stage Training
To achieve a balance between image generation and understanding capabilities, our training process is carried out in two stages. The first stage utilizes large-scale, mid-quality data (as illustrated in fig. 2) to enhance the diversity of the model’s data distribution (Hu et al., 2024). By initially modeling this diverse data, we strengthen the model’s understanding capabilities. In the second stage, we employ a smaller volume of high-quality data to further improve the image generation capacity and refine the model’s comprehension of images.
We observe that when the training dataset is excessively large, preventing the model from iterating over each image hundreds of times, maintaining an image mask ratio within the range of [0.7, 1], as done in MAR, hinders the model’s ability to generate coherent image tokens. Consequently, in the second stage, we adjust the mask ratio to (0, 1]. The example is showed in Appendix A.5. The probability density curve of the two mask ratios are shown on the right side of fig. 2. Additionally, due to the small size of the diffusion mlp and the limited amount of high-quality data, we increased the number of timestep samples for the given $z$ to enhance learning efficiency. To further mitigate performance fluctuations in the model, we employed an exponential moving average (EMA) method with a momentum of 0.9999 in both stages.
3.3.4 Inference Strategy
Classifier-Free Guidance
During the inference stage, the model performs both text-based and non-text-based image generation tasks. The provided conditions are represented as $z_{c}$ and $z_{u}$, and the predicted $v$ is given by: $v=v_{\theta}(x^{(t)}|t,z_{u})+\omega*(v_{\theta}(x^{(t)}|t,z_{c})-v_{\theta}(x^{(t)}|t,z_{u}))$ (which has the same effect as $\epsilon=\epsilon_{u}+\omega*(\epsilon_{c}-\epsilon_{u})$ , with the mathematical derivation process detailed in the Appendix A.4), where $\omega$ denotes the guidance scale. Our method has been experimentally validated to support a large guidance scale, as compared to the noise prediction training approach. We hypothesize that this is due to the strong conditioning provided by $z$, which, coupled with the auto-regressive (AR) process, necessitates reducing accumulated error. Thus, each token must be generated as accurately as possible, a requirement that can be fulfilled with a large guidance scale.
Mixed-Precision Inference
To enhance numerical stability during image generation, particularly when using small step sizes in the DDPM sampling process, we perform the model’s forward pass in bfloat16 (matching the training precision) but cast the output to float32 before DDPM sampling. This mitigates potential rounding error without a significant computational overhead, therefore improving the sampling accuracy efficiently.
4 Experiment
4.1 Dataset
We utilized the Capfusion-120M dataset for the image expert pretraining stage. This dataset is publicly accessible and comprises an extensive collection of web-based image-text pairs, designed to optimize noisy captions (Yu et al., 2023). In an effort to further improve the quality of the content generated, we executed a random sampling of 20M data points from the Capfusion dataset during our image expert fine-tuning stage. This was supplemented with a high-quality mixed dataset that included ImageNet-1K-1.2M (Deng et al., 2009), CC12M (Changpinyo et al., 2021), and LAION-aesthetics-12M [^1] https://huggingface.co/datasets/dclure/laion-aesthetics-12m-umap. We employ the open-source InternVL2-8B for recaptioning the CC12M and laion-aesthetics-12m datasets in English. Following LLaVA-v1.5 (Liu et al., 2023a), we use LLaVA-v1.5-mix-665K data for instruction tuning before each performance test for visual understanding.
4.2 Implementation Details
By default, we utilized the AdamW optimizer with betas (0.9, 0.95). The weight decay was consistently maintained in proportion to the learning rate. During the first training stage, a learning rate of 5e-5 was employed for a total of 4 epochs, with an initial warm-up phase of 0.5 epochs, followed by maintaining the learning rate at 5e-5. In the second stage, the 0.5B model maintained a 5e-5 learning rate and a 0.5 epoch warm-up, with training lasting for 3 epochs. For the larger 7B model, we initially applied a 2e-6 learning rate for 1.5 epochs, before transitioning to a 1e-6 learning rate for an additional 1.5 epochs. Furthermore, during the first stage, the total batch size for the smaller 0.5B model was 2496, while it was 1152 for the larger 7B model. In the second stage, the total batch size for the smaller 0.5B model was 768, and for the larger 7B model, it was 480. Notably, in both stages, only the PLoRA portion of the LLM parameters is trainable, and a consistent image resolution of 256x256 was used throughout. Our models utilize bfloat16 during training and inference. For DDPM sampling, we convert the model output to float32 for higher precision.
4.3 Comparison with Other Systems
Table 1: Comparison on visual understanding benchmarks. MMAR surpasses other joint image-text probabilistic models by a large margin, even with a small resolution of 256x256, approaching the performance of traditional MLLMs like LLaVA, which employ pretrained CLIP vision encoder.
Method LLM V-Token Res. MMB MMEP POPE SEED MM-Vet AVE@18Und.
LLaVA-1.5 Vicuna-1.5-7B CLIP 336 64.3 1510.7 85.9 58.6 31.1 47.08
EMU-2 LLaMA-13B CLIP 448 62.8 48.5
SEED-X LLaMA-13B CLIP dynamic 75.4 1435.7 84.2
DreamLLM Vicuna-7B CLIP 224 58.2 36.6
Chameleon-7B 7B from scratch vq-vae 512 13.32 125.8 30.86 34.61 7.34 18.34
Transfusion* Qwen-2-0.5B vae 256 29.47 594.3 66.90 42.40 13.90 28.26
Show-o Phi-1.5B CLIP 336 42.44 1182.7 84.50 51.61 20.87 33.06
VILA-U LLaMA-2-7B vq-vae 256 1336.2 83.9 56.3 27.7
MMAR-0.5B Qwen-2-0.5B vae 256 48.45 882.1 70.74 55.70 18.49 34.56
MMAR-7B Qwen-2-7B vae 256 66.32 1393.9 83.02 64.52 27.80 46.52
Visual Understanding. As depicted in table 1, we thoroughly gauge MMAR’s performance in visual understanding, employing VLMEvalKit (Duan et al., 2024) to perform extensive evaluations on prevalent visual understanding benchmarks, encompassing a total of 18 such assessments (average score denoted by “AVE@18Und.” in table 1, for details see Appendix A.7.) including MMB (Liu et al., 2023b), MME (Fu et al., 2023), POPE (Li et al., 2023c), SEED (Li et al., 2023a), MM-Vet (Yu et al., 2024), among others. Our method outperforms other joint image-text probabilistic models by a large margin, including Chameleon-7B (Team, 2024), Show-o (Xie et al., 2024), VILA-U (Wu et al., 2024) and our re-implemented version of Transfusion (denoted by “Transfusion*”), approaching the performance of traditional MLLMs like LLaVA, which employ pretrained CLIP vision encoder. Even without using any pre-trained CLIP or diffusion models and with small resolution of 256 $\times$ 256, MMAR-7B presents comparable or even better performance when compared to methods using pre-trained clip and diffusion models, including EMU-2 (Sun et al., 2024c), SEED-X (Ge et al., 2024), and DreamLLM (Dong et al., 2024a).
Visual Generation. We showcase the zero-shot FID of MMAR, evaluated on the MSCOCO 30k dataset, in Table 2. Our model’s performance is discernibly on par with current robust generative models, a noteworthy achievement for MMAR given its minimal training epochs. Repeated exposure to identical data can substantially enhance a model’s generative quality, as exemplified by MAR, which trains for 400 epochs on the 1.2M ImageNet dataset, and Show-o, which undergoes approximately 49 epochs of training on the 35M dataset.
Table 2: Comparison on MSCOCO Dataset
[…]
Gen. Only DALL-E 12B 250M 27.50
LDM 1.4B 400M 12.64
DALL-E2 6.5B 650M 10.39
Imagegen 3B 5000M+ 6.61
Und. and Gen. w/ pre-trained Diff. CoDI - 400M 11.26
SEED-X 17B - 12.68
DreamLLM 7B - 8.76
Joint Prob. Models Show-o 1.3B 35M 9.24
Chanmeleon 7B - 29.6
Transfusion  [^2] Both the Transfusion and Chanmeleon results are referenced from Table 3 in the paper ‘Transfusion: Predict the Next Token and Diffuse Images with One Multi-Modal Model.’ 7B - 16.8
MMAR-0.5B 0.5B 145.2M 36.6
MMAR-7B 7B 145.2M 17.1
However, to bolster the model’s comprehension capabilities, exposure to diverse data is crucial rather than simply reiterating the same data. Consequently, we restrict our training to just 3 epochs for the second stage high-quality dataset. Despite these constraints, our model maintains competitive performance in generation, further substantiating the efficacy of our image-text joint modeling approach. The example images generated by our model are shown in the Appendix A.6.
Scaling up with Model Size. From Table 1 and 2, we see that MMAR can scale up from 0.5B to at least 7B model size, with significant improvement of visual understanding and generation capability.
4.4 Ablation Study
We begin by evaluating the impact of our chosen diffusion parameterization method. Table 3 demonstrates that switching to the more common $\epsilon$-prediction leads to a significant decrease in both visual understanding and generation quality, confirming the effectiveness of our optimal diffusion parameterization. To evaluate the efficacy of various image-text joint modeling techniques, we devise two distinct versions based on the MMAR framework: one employing LDM-VQ-16 tokenizer (Rombach et al., 2022) for discrete token modeling, and the other utilizing Transformer for diffusion modeling (refer to Transfusion). Comprehensive implementation details are provided in the Appendix A.1. Our test results are presented in Table 3. The Transformer-based diffusion modeling version considerably underperforms the other two approaches in both understanding and generation aspects. This is attributed to the substantial loss of image information when jointly modeling image-text and operating with limited training epochs. In contrast, our method consistently delivers superior results.
Table 3: Ablation study on MMAR
[…]
MMAR-0.5B(Full method) 48.45 882.1 70.74 55.70 18.49 34.56 36.6
w/ $\epsilon$-pred. 45.53 880.7 71.14 53.72 17.98 32.21 61.53
w/ VQ 37.54 618.2 66.98 44.93 14.45 29.70 66.26
transfusion-like 29.47 594.3 66.90 42.40 13.90 28.26 95.38
4.5 Analysis
Impact of v-prediction
To delve deeper into the disparities between v-prediction and $\epsilon$-prediction for diffusion MLP, we independently collect the statistics of the MSE loss of $v^{(t)}$ values at various time steps $t$ throughout the training process for both methods, as illustrated in fig. 3 (A). Furthermore, to more effectively discern the loss discrepancies between the two techniques, we subtract the v-prediction curve from the $\epsilon$-prediction curve, yielding the yellow curve. For reference, we plot the theoretically predicted numerical error of the $\epsilon$-prediction model (see Appendix A.3) as the red curve. The graph reveals that the loss of $\epsilon$-prediction model is consistently higher than v-prediction model. Especially, when $t>900$, curve “$\epsilon-v$” exhibits a significant spike towards infinity. This aligns with the behavior of the theoretical numerical error, confirming the non-negligible numerical error effect in low-precision training. The gap between yellow and red curve indicates that apart from its direct effect, numerical error also introduces optimization difficulty, hindering the loss convergence.
Impact of CFG Scaling
We select models from the second and fourth epochs of the first stage as starting points for the second stage, train them for 3 epochs, and then test the MSCOCO FID-30K under varying CFG intensities. As shown in fig. 3 (B), our method achieves better FID scores as the CFG scale increases from 1 to 10. It is worth noting that most probabilistic generative models typically have a CFG scale between 1.5 and 5. Additionally, it is observed that a longer training duration in the first stage consistently results in better generation outcomes at all CFG scales.
Scaling up with More Training Data
As illustrated in fig. 3 (C), we evaluate the average performance of MMAR-0.5B on 18 visual understanding benchmarks, following SFT training applied to the checkpoints generated during its training process. The blue background in the figure denotes the Image Expert Pretraining stage, while the green background signifies the Image Expert Fine-tuning stage. The curve reveals that, with an increasing number of training steps, i.e. more training data, the visual understanding performance of MMAR-0.5B consistently improves without reaching saturation. This finding highlights the exceptional scale-up capability of our MMAR model.
[…]
Figure 3: The impact of v-prediction, CFG scaling, and training steps.
4.6 Limitation
Our method still requires further optimization in terms of image generation speed. Although the diffusion mlp is not involved in image understanding tasks, when applied to image generation tasks, we are compelled to use a 256-step token generation followed by a 100-step diffusion denoising process to ensure the quality of the generated images. This results in a generation time of nearly three minutes for a single 256x256 image using MMAR-7B. While it is possible to generate multiple images simultaneously by increasing the batch size, this does not fundamentally resolve the issue of prolonged generation time. We plan to address this in future work.
5 Conclusion
This paper proposes MMAR, a novel multimodal auto-regressive probabilistic modeling framework based on continuous image representations. It employs a standalone diffusion mlp at the image token level on top of pre-trained LLMs to facilitate theoretical lossless joint image-text probabilistic modeling. In practice, the low precision training of LLMs poses an non-negligible numerical error term to diffusion loss, causing optimization difficulty. This was addressed by deriving an optimal diffusion model parameterization. To balance the understanding and generation ability, a two-stage training strategy is introduced. During inference time, MMAR can tolerant extremely large CFG scale to generate high quality images. MMAR significantly demonstrates scaling-up laws with more data and larger model size. Extensive evaluations are conducted on 18 image understanding benchmarks, revealing that MMAR is the first joint image-text modeling framework that approaches comparable performance with traditional MLLMs that employ pretrained CLIP vision encoder, marking a significant step toward lossless joint probabilistic modeling of images and text.
Appendix A Appendix
A.1 Additional Implementation Details
VQ
Based on the MMAR-0.5B framework, we replace the Image Tokenizer from KL-16 to VQ-16. The image codes extracted using VQ-16 are then passed through a projector to increase the channel size to match the LLM’s hidden size. Subsequently, we add a decoding Linear layer, which takes the hidden states of the LLM’s output image portion as input and maps them to the image codebook. The Cross Entropy loss is then calculated between these mapped values and the actual VQ codes.
Transfusion
Following the theoretical ideas presented in the Transfusion paper, we adopt a simple linear mapping. After extracting image tokens using KL-16, if the task is image first, we add noise within a 500-time step to the image tokens. Otherwise, we add noise within a 1000-step. After the linear mapping, we add the image token a learnable time embedding corresponding to the time step as input to the LLM. We also maintain the bidirectional attention mechanism.
After passing through the LLM, we first map the output back to the original token channel count using a linear layer and compute the MSE loss for the predicted noise. During generation, we treat LLM as a denoised model, with the condition being the concatenation of the text and the image tokens to be generated. We adopt the learning approach of Transfusion but conduct experiments based on our training tasks and stage divisions.
Projector
In order to accomplish channel alignment, we introduced two Projectors designed to scale the channels. Both Projectors consist of a simple linear layer and multiple blocks composed of activation layers and linear layers. The PostProjector comprises one block, whereas the VisProjector contains two blocks.
A.2 Minimizing the Numerical Error in Diffusion Models
To make our discussion clearer, we switch the diffusion noise schedule into an angular form as follows:
$\begin{cases}\sin\phi_{t}=\sqrt{1-\bar{\alpha}_{t}},\\ \cos\phi_{t}=\sqrt{\bar{\alpha}_{t}}.\end{cases}$ (9)
In this way, the forward diffusion process can be written as follows:
$x^{(t)}=\sqrt{\bar{\alpha}_{t}}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon=\cos\phi_{t}x+\sin\phi_{t}\epsilon,$ (10)
where $x^{(t)}$, $x$ and $\epsilon$ are noised image latent, original image latent and gaussian noise, respectively.
Our goal is to minimize the numerical error term in the DDIM sampling process. However, the form of DDIM sampling process is different under different parameterization method of the diffusion model. Therefore, we need to first define a general form to represent the diffusion model parameterization.
We consider the diffusion model output $u_{\theta}^{(t)}$ predict a linear combination of data $x$ and noise $\epsilon$, i.e. $u^{(t)}=a_{t}x+b_{t}\epsilon$. Note that the coefficients can vary according to the diffusion time step $t$. Further re-writing the coefficients in the angular form gives:
$u^{(t)}=r_{t}\cos\psi_{t}x+r_{t}\sin\psi_{t}\epsilon,$ (11)
where $r_{t}=\sqrt{a_{t}^{2}+b_{t}^{2}}$ represents the scale of $u^{(t)}$. $\cos\psi_{t}$ and $\sin\psi_{t}$ balance the proportion of $x$ and $\epsilon$. Combining eq.10 and eq.11, we can in turn represent $x$ and $\epsilon$ with $u^{(t)}$ and $x^{(t)}$:
$\begin{cases}x=\frac{\sin\psi_{t}x^{(t)}-\sin\phi_{t}u^{(t)}/r_{t}}{\cos\phi_{t}\sin\psi_{t}-\cos\psi_{t}\sin\phi_{t}}=\frac{\sin\psi_{t}x^{(t)}-\sin\phi_{t}u^{(t)}/r_{t}}{\sin(\psi_{t}-\phi_{t})},\\ \epsilon=\frac{\cos\psi_{t}x^{(t)}-\cos\phi_{t}u^{(t)}/r_{t}}{\sin\phi_{t}\cos\psi_{t}-\sin\psi_{t}\cos\phi_{t}}=-\frac{\cos\psi_{t}x^{(t)}-\cos\phi_{t}u^{(t)}/r_{t}}{\sin(\psi_{t}-\phi_{t})}.\end{cases}$ (12)
Now we consider the general form of DDIM sampling step (Song et al., 2021):
$x^{(t-1)}=\cos\phi_{t-1}\hat{x}_{\theta}(x^{(t)})+\sin\phi_{t-1}\hat{\epsilon}_{\theta}(x^{(t)}),$ (13)
where $\hat{x}_{\theta}(x^{(t)})$ and $\hat{\epsilon}_{\theta}(x^{(t)})$ are the estimated image latent and noise, respectively.
Note that by using eq.12, both of $\hat{x}_{\theta}(x^{(t)})$ and $\hat{\epsilon}_{\theta}(x^{(t)})$ can be derived from the noisy image latent $x^{(t)}$ and the diffusion model output $u_{\theta}^{(t)}$. Therefore, we can further represent $x^{(t-1)}$ in the following form:
$x^{(t-1)}$ $=\cos\phi_{t-1}\frac{\sin\psi_{t}x^{(t)}-\sin\phi_{t}u_{\theta}^{(t)}/r_{t}}{\sin(\psi_{t}-\phi_{t})}-\sin\phi_{t-1}\frac{\cos\psi_{t}x^{(t)}-\cos\phi_{t}u_{\theta}^{(t)}/r_{t}}{\sin(\psi_{t}-\phi_{t})}$
$=\frac{\sin(\phi_{t-1}-\phi_{t})u_{\theta}^{(t)}/r_{t}-\sin(\phi_{t-1}-\psi_{t})x^{(t)}}{\sin(\psi_{t}-\phi_{t})}.$ (14)
Eq.14 represents the general form of DDIM sampling step under any kind of diffusion model parameterization in the form of eq.11. To help understanding, we further present the geometric meaning of eq.14. As shown in fig.4, term $x^{(t-1)},x^{(t)}$, and $u_{\theta}^{(t)}/r_{t}$ all locate on the unit circle in the $x-\epsilon$ plain. We find that eq.14 can be interpreted as projecting $x^{(t-1)}$ onto the $(x^{(t)},\frac{u_{\theta}^{(t)}}{r_{t}})$ coordinate system. We illustrate this projection by adding auxiliary line $AB$ and $AC$. By solving the sine law of $\triangle OBA$ given $OA=1$, we get:
$\begin{cases}OB=\frac{\sin(\Delta\phi)}{\sin(\psi_{t}-\phi_{t})}\\ BA=\frac{-\sin(\phi_{t-1}-\psi_{t})}{\sin(\psi_{t}-\phi_{t})}\end{cases}$ (15)
By representing $x^{(t-1)}=OB\cdot u_{\theta}^{(t)}/r_{t}+AB\cdot x^{(t)}$, we get:
$x^{(t-1)}=\frac{\sin(\Delta\phi)}{\sin(\psi_{t}-\phi_{t})}u_{\theta}^{(t)}/r_{t}-\frac{\sin(\phi_{t-1}-\psi_{t})}{\sin(\psi_{t}-\phi_{t})}x^{(t)},$ (16)
which aligns with eq.14 given that $\Delta\phi=\phi_{t-1}-\phi_{t}$.
Figure 4: Geometric interpretation of a DDIM sampling step under arbitrary diffusion model parameterization.
Now, we take the numerical error into consideration by multiplying the model output by a factor $1+\delta$, where $\delta$ represents the relative error:
$\tilde{x}^{(t-1)}=\frac{\sin(\phi_{t-1}-\phi_{t})(1+\delta)u_{\theta}^{(t)}/r_{t}-\sin(\phi_{t-1}-\psi_{t})x^{(t)}}{\sin(\psi_{t}-\phi_{t})}.$ (17)
Further, we can isolate the numerical error term from the ideal DDIM sampling step:
$\tilde{x}^{(t-1)}=x^{(t-1)}+\sin(\phi_{t-1}-\phi_{t})\frac{u_{\theta}^{(t)}/r_{t}}{\sin(\psi_{t}-\phi_{t})}\delta.$ (18)
From eq.18, we conclude that the numerical error of a DDIM sampling step is determined by four factors, namely, the step size $\Delta\phi=\phi_{t-1}-\phi_{t}$, the normalized model output $u_{\theta}^{(t)}/r_{t}$, the relative error of the data type $\delta$, and $\sin(\psi_{t}-\phi_{t})$, which is decided by the parameterization of the diffusion model.
Notably, not all these four factors are useful to achieve the goal of minimizing the numerical error. For example, tuning down the step size only decreases the numerical error of each step. As a result, the total step number of DDIM sampling is increased proportionally, which cancels out the effect of error reduction of each single step. The factor $u_{\theta}^{(t)}/r_{t}$ is not adjustable since it constantly has a unit standard deviation. This can be verified by the following calculation:
$\mathbb{E}[(u^{(t)}/r_{t})^{2}]=\mathbb{E}[(\cos\psi_{t}x+\sin\psi_{t}\epsilon)^{2}]=\cos^{2}\psi_{t}\mathbb{E}[x^{2}]+\sin^{2}\psi_{t}.$ (19)
In common practice, image tokens $x$ are normalized into unit standard deviation. Therefore, $\mathbb{E}[(u^{(t)}/r_{t})^{2}]=\cos^{2}\psi_{t}+\sin^{2}\psi_{t}=1$.
If we decide to scale up our model, it is better to leverage the pre-trained LLMs as well as the highly efficient training infrastructure that is specifically optimized for LLMs. This makes bfloat16 almost the only choice. As a result, the relative error $\delta$ is fixed to $1/128$.
Now, our only choice is to adjust the diffusion model parameterization method, so that $|\sin(\psi_{t}-\phi_{t})|$ is maximized. A simple solution is to set $\psi_{t}-\phi_{t}=\pi/2$, resulting in the following parameterization:
$u^{(t)}=r_{t}\cos(\phi_{t}+\pi/2)x+r_{t}\sin(\phi_{t}+\pi/2)\epsilon=r_{t}(\cos\phi_{t}\epsilon-\sin\phi_{t}x).$ (20)
Note that $r_{t}$ is still undetermined, which reflects the scale of $u^{(t)}$. From the analysis above, $r_{t}$ does not affect the numerical error term, since it is canceled out by the normalization of the model output, as seen in the factor $u_{\theta}^{(t)}/r_{t}$. Therefore, $r_{t}$ can be chosen freely, or based on other considerations. We consider that the smooth optimization of a neural network often requires the activation and output not too large or small. Therefore, we require a unit standard deviation for $u^{(t)}$, making $r_{t}=1$ constantly.
The final parameterization of our diffusion model is as follows:
$u^{(t)}=\cos\phi_{t}\epsilon-\sin\phi_{t}x.$ (21)
We notice that this parameterization is coincidentally the “v-prediction” parameterization (Salimans & Ho, 2022). Note that, however, “v-prediction” is initially proposed for the efficient distillation of diffusion models, rather than reducing the numerical error of diffusion models. To the best of our knowledge, our work is the first to derive “v-prediction” parameterization from the first principle of minimizing the numerical error in diffusion models.
A.3 Deriving Theoretical Numerical Error for $\epsilon$-Prediction Models
The $\epsilon$-prediction parameterization corresponds to $\psi_{t}=\frac{\pi}{2}$ in the angular parameterization form given by eq.11. Substituting $\psi_{t}=\frac{\pi}{2}$ and $u^{(t)}_{\theta}/r_{t}=\epsilon_{\theta}$ into eq.18, we get:
$\tilde{x}^{(t-1)}=x^{(t-1)}+\sin(\phi_{t-1}-\phi_{t})\frac{\epsilon_{\theta}}{\cos(\phi_{t})}\delta.$ (22)
Further, we cancel out the step size factor $\sin(\phi_{t-1}-\phi_{t})$ within the numerical error term, only focusing on “the numerical error introduced per unit DDIM step”:
$e^{(t)}=\frac{\epsilon_{\theta}}{\cos\phi_{t}}\delta.$ (23)
Next, we will show that $e^{(t)}$ can also be interpreted as the equivalent v-prediction numerical error for an $\epsilon$-prediction model.
For an $\epsilon$-prediction model, $u_{\theta}^{(t)}=\epsilon_{\theta}$. In order to calculate the equivalent $v^{(t)}_{\theta}$ value, we need to represent $v^{(t)}_{\theta}$ with the predicted $\epsilon_{\theta}$ and the known $x^{(t)}$, which is calculated as follows:
$v^{(t)}_{\theta}=\cos\phi_{t}\epsilon_{\theta}-\sin\phi_{t}\hat{x}_{\theta}(x^{(t)})=\cos\phi_{t}\epsilon_{\theta}-\sin\phi_{t}\frac{x^{(t)}-\sin\phi_{t}\epsilon_{\theta}}{\cos\phi_{t}}=\frac{\epsilon_{\theta}}{\cos\phi_{t}}-\tan\phi_{t}x^{(t)}.$ (24)
Considering the numerical error, we get:
$\tilde{v}^{(t)}_{\theta}=\frac{\epsilon_{\theta}(1+\delta)}{\cos\phi_{t}}-\tan\phi_{t}x^{(t)}=v^{(t)}_{\theta}+\frac{\epsilon_{\theta}}{\cos\phi_{t}}\delta.$ (25)
Note that the numerical error term in the above equation is exactly $e^{(t)}$, proving that $e^{(t)}$ can be interpreted as the equivalent v-prediction numerical error for an $\epsilon$-prediction model.
Taking numerical error effect into the v-prediction-based diffusion loss, we get:
$\mathbb{E}[(v^{(t)}-\tilde{v}^{(t)}_{\theta})^{2}]=\mathbb{E}[(v^{(t)}-v^{(t)}_{\theta}-e^{(t)})^{2}]=\mathbb{E}[(v^{(t)}-v^{(t)}_{\theta})^{2}]-2\mathbb{E}[(v^{(t)}-v^{(t)}_{\theta})e^{(t)}]+\mathbb{E}[(e^{(t)})^{2}].$ (26)
Due to the fact that numerical error $e^{(t)}$ is independent from the training loss and that the expectation of $e^{(t)}$ is 0, we get $\mathbb{E}[(v^{(t)}-v^{(t)}_{\theta})e^{(t)}]=0$. Therefore, the only numerical error term is $\mathbb{E}[(e^{(t)})^{2}]$. Given that the standard deviation of $\epsilon_{\theta}$ is 1, and considering that we use bfloat16 as training data type, which means $\delta=1/128$, we get
$\mathbb{E}[(e^{(t)})^{2}]=1/(128\cos(\phi_{t}))^{2}=1/(128^{2}\bar{\alpha_{t}}).$ (27)
This is the theoretical numerical error of the v-prediction diffusion loss for an $\epsilon$-prediction model.
A.4 CFG With $v$-prediction
From Equation $v^{(t)}_{i}=\sqrt{\bar{\alpha}_{t}}\mathbf{\epsilon}-\sqrt{1-\bar{\alpha}_{t}}x_{i}$, we can derive the following equation.
$\epsilon=\sqrt{1-\bar{\alpha}^{(t)}}x^{(t)}+\sqrt{\bar{\alpha}^{(t)}}v$ (28)
For the CFG of $\epsilon$, it can be simplified as follows.
ϵ=italic-$\epsilon=$ $\epsilon_{u}+\omega(\epsilon_{c}-\epsilon_{u})$
$=$ $\sqrt{1-\bar{\alpha}^{(t)}}x^{(t)}+\sqrt{\bar{\alpha}^{(t)}}v_{u}+\omega\sqrt{\bar{\alpha}^{(t)}}(v_{c}-v_{u})$
$=$ $\sqrt{1-\bar{\alpha}^{(t)}}x^{(t)}+\sqrt{\bar{\alpha}^{(t)}}(v_{u}+\omega(v_{c}-v_{u}))$ (28)
Ultimately, we obtain $v=v_{u}+\omega(v_{c}-v_{u})$. The CFG of $v$ and $\epsilon$ are equivalent.
A.5 Examples:The effect of the stage 2 training
Figure 5: The impact of the second training stage on image generation capability.
A.6 Examples:Image Generation
Figure 6: Generated images from MMAR-7B
A.7 Detailed Visual Understanding Evaluation Results
A total of 18 visual understanding benchmarks from VLMEvalKit (Duan et al., 2024) are used to evaluate MMAR models comprehensively. The evaluation is also conducted on the existing joint image-text probabilistic models using the publicly available checkpoints [^3] https://huggingface.co/facebook/chameleon-7b [^4] https://huggingface.co/showlab/show-o-w-clip-vit. The detailed evaluation results are shown in table 4. All scores have been scaled to a range of 0 to 100 except that we show the original score of MME benchmarks. The average score is calculated on the normalized score of all the benchmarks including MME.
Table 4: Detailed visual understanding evaluation results.
Benchmark Chameleon-7B Transfusion* Show-o MMAR-0.5B MMAR-0.5B w/ $\epsilon$-pred. MMAR-0.5B w/ VQ MMAR-7B
AI2D 34.81 40.22 32.48 43.43 41.90 41.90 62.63
ChartQA 3.84 9.56 11.32 10.20 10.36 9.36 12.72
DocVQA 1.51 6.72 18.24 7.62 6.77 6.79 10.28
Hallu.Bench 39.01 41.54 40.90 42.80 41.11 41.54 50.47
MathVista 21.90 22.60 23.20 21.60 23.10 22.90 32.10
MMBenchCN 10.14 27.23 0.52 43.99 38.83 31.87 66.49
MMBenchEN 13.32 29.47 42.44 48.45 45.53 37.54 66.32
MMEP 125.8 594.3 1182.7 882.1 880.7 618.2 1393.9
MMEC 33.9 206.1 225.0 256.8 232.1 273.2 320.0
MMMU 24.00 29.33 26.44 29.33 25.33 29.67 42.78
MMStar 20.47 28.13 32.00 32.13 31.07 28.07 40.00
MMVet 7.34 13.90 20.87 18.49 17.98 14.45 27.80
OCRBench 0.50 2.30 15.20 18.70 7.10 2.10 23.10
POPE 30.86 66.90 84.50 70.74 71.14 66.98 83.02
RealWorldQA 27.06 36.99 27.97 38.30 35.16 36.60 52.42
ScienceQA 44.83 45.92 41.82 47.54 45.21 45.35 71.05
SEEDBench 34.61 42.40 51.61 55.70 53.72 44.93 64.52
TextVQA 5.43 9.94 38.35 16.77 12.40 9.46 22.05
Average 18.34 28.26 33.06 34.56 32.21 29.70 46.52
A.8 Analysis on the Visual Understanding Ability of Transfusion-Like Methods
In this section, we will quantitatively analyze the incomplete information usage problem in transfusion-like methods.
The experiments are conducted using our re-implemented transfusion model. The model is evaluated on MMBench using different noise-level image input. Specifically, we take the diffusion time-step of 0, 125, 250, 375, 500, and 750 for testing. Among the MMBench tasks, we focus on tasks whose performance is significantly affected by noise-levels. Fig. 7 depict the performance trend of these tasks across different noise levels.
Figure 7: The trend of MMBench tasks’ performance with respect to input image noise levels. (A): The best performance of most tasks are achieved at a non-zero noise level. (B): The object localization performance consistently gets higher as the noise level grows up.
Intuitively, increasing noise levels should lead to a consistent performance decline across all visual understanding tasks due to information loss. However, our findings reveal a more nuanced relationship. We find that the best performance of most visual understanding tasks are achieved at a non-zero noise level. Especially, the object localization task, which is more focused on the image layout information, consistently performs better as the noise level grows higher. This suggests that transfusion-like methods, when limited to a single noise level input, fail to fully exploit the information of the input image. Instead, visual information appears to be distributed across different noise levels within the model, with the input noise level dictating which information is ultimately utilized.

### v3

@@
 Recent advancements in multi-modal large language models have propelled the development of joint probabilistic models capable of both image understanding and generation.
-However, we have identified that recent methods inevitably suffer from loss of image information during understanding task, due to either image discretization or diffusion denoising steps.
+However, we have identified that recent methods suffer from loss of image information during understanding task, due to either image discretization or diffusion denoising steps.
 To address this issue, we propose a novel Multi-Modal Auto-Regressive (MMAR) probabilistic modeling framework.
-Unlike discretization line of method, MMAR takes in continuous-valued image tokens to avoid information loss.
+Unlike discretization line of method, MMAR takes in continuous-valued image tokens to avoid information loss in an efficient way.
 Differing from diffusion-based approaches, we disentangle the diffusion process from auto-regressive backbone model by employing a light-weight diffusion head on top each auto-regressed image patch embedding.
@@
 To successfully train our method, we also propose a theoretically proven technique that addresses the numerical stability issue and a training strategy that balances the generation and understanding task goals.
-Through extensive evaluations on 18 image understanding benchmarks, MMAR demonstrates much more superior performance than other joint multi-modal models, matching the method that employs pretrained CLIP vision encoder, meanwhile being able to generate high quality images at the same time.
-We also showed that our method is scalable with larger data and model size.
+Extensive evaluations on 18 image understanding benchmarks show that MMAR significantly outperforms most of the existing joint multi-modal models, surpassing the method that employs pre-trained CLIP vision encoder.
+Meanwhile, MMAR is able to generate high quality images.
+We also show that our method is scalable with larger data and model size.
 Code will be available at https://github.com/ydcUstc/MMAR.
@@
 Over the past few years, extensive research in the field of multimodal intelligence has catalyzed the accelerated advancement of foundational models for both image understanding and image generation.
-Within the realm of image understanding, multimodal large language models (MLLM), exemplified by LLaVA (Liu et al., 2023a), have exhibited human-like capabilities in open-domain image comprehension.
-In the domain of image generation, techniques rooted in generative probabilistic models, such as Denoising Diffusion Probabilistic Models (DDPM) (Ho et al., 2020) and auto-regressive (AR) models (Chen et al., 2020), have also garnered significant success.
+Within the realm of image understanding, multimodal large language models (MLLM), exemplified by LLaVA [33], have exhibited human-like capabilities in open-domain image comprehension.
+In the domain of image generation, techniques rooted in generative probabilistic models, such as Denoising Diffusion Probabilistic Models (DDPM) [21] and auto-regressive (AR) models [6], have also garnered significant success.
 Essentially, these two lines of research correspond to modeling the conditional probability, i.e., $P(T|I)$ and $P(I|T)$, where $T$ and $I$ corresponds to text and image, respectively.
@@
 This brings us to an intriguing question: Could a joint probabilistic model serve as a natural unifying framework for both image understanding and image generation tasks?
-Given that the most advanced image understanding (Chen et al., 2024) and generation (Esser et al., 2024) models rely on the language priors $p(T)$ of pre-trained large language models (LLMs), the most straight-forward approach for joint image-text probabilistic modeling is to convert images into discrete tokens similar to text.
-This way, images are treated as a form of language and integrated into the auto-regressive modeling of text, as seen in models like MARS (He et al., 2024), LlamaGen (Sun et al., 2024a), and Chameleon (Team, 2024).
+Given that the most advanced image understanding [8] and generation [14] models rely on the language priors $p(T)$ of pre-trained large language models (LLMs), the most straight-forward approach for joint image-text probabilistic modeling is to convert images into discrete tokens similar to text.
+This way, images are treated as a form of language and integrated into the auto-regressive modeling of text, as seen in models like MARS [19], LlamaGen [53], and Chameleon [57].
 This method leverages the powerful text modeling capabilities of various open-source pre-trained LLMs, along with their highly optimized training and inference frameworks.
-However, discretization on the continuous image space introduces an information bottleneck, inevitably leading to a loss of image details and reducing the model’s information capacity.
-This limitation is quantitatively evident in the image understanding performance: existing methods based on image token discretization fall short when compared to the LLaVA model, which utilizes continuous CLIP representations (Xie et al., 2024).While there are theoretically feasible methods to mitigate this information bottleneck, such as significantly increasing the number of tokens or the size of the VQ codebook, these approaches would substantially increase the training difficulty for both LLMs and VQ-VAE models.
-Recently, several attempts have been made to address the continuous nature of images by combining image diffusion transformers and text auto-regression within a unified transformer architecture (Zhou et al., 2024; Zhao et al., 2024).
-Although this approach, exemplified by Transfusion (Zhou et al., 2024) and MonoFormer, demonstrated superior performance compared to discrete token auto-regressive methods for images, it is crucial to acknowledge that the inherent nature of image diffusion prevents this approach from leveraging complete image modeling capability when it comes to image understanding tasks, as shown in fig.
-1-(B).
-The root cause is that image diffusion will encode different parts of image information into different noise levels (Ho et al., 2020).
-For example, at low noise level, the model tends to solve an image enhancement task, while at high noise level, the model is only suitable for extracting rough image morphology and layout.
-Intuitively, image-to-text tasks need to take clean images as input to maintain complete visual information.
-However, due to the nature of diffusion modeling, this will only invoke the hidden representations that are related to the image enhancement task according to its low noise level, resulting in insignificant help for image-to-text tasks.
+However, the code-book size strictly restricts the information capacity per token, leading to a loss of image details and reducing the model’s information capacity when the image token number is small.
+This limitation is quantitatively evident in the image understanding performance: most of the existing methods based on image token discretization [57, 62] fall short when compared to the LLaVA model, which utilizes continuous CLIP representations.
+Although some approaches can mitigate this bottleneck, such as increasing the number of tokens (4096 tokens per 512x512 image for EMU3 [60]) or increasing the size of the VQ code-book, these approaches would substantially increase the training cost for both LLMs and VQ-VAE models.
+In contrast, without the restriction of code-book size, continuous image tokenizers can compress images much more efficiently.
+Recent research [4] even shows that it is possible to compress a 128x128 image patch into a single continuous image token without significant detail loss, resulting in only 16 tokens per 512x512 image.
+This indicates huge potential of using continuous image tokens instead of discrete ones.
+Figure 1: The trend of MMBench tasks’ performance of a transfusion-like model with respect to input image noise levels.
+(A): The best performance of most tasks are achieved at a non-zero noise level.
+(B): The object localization performance consistently gets higher as the noise level grows up.
+Recent efforts on leveraging continuous image tokens include transfusion [71] and MonoFormer [70].
+They combine image diffusion transformers and text auto-regression within a unified transformer architecture, and demonstrates superior performance compared to image discretization approaches.
+However, when we reproduce and analyze this approach, we find it has difficulty utilizing complete image modeling capability for image understanding tasks when only the clean image is input to the model.
+As shown in Fig.
+1, the best performance of different image understanding tasks are achieved at different non-zero noise levels.
+Especially, the object localization task even prefers very noisy inputs.
+This counter-intuitive phenomenon reveals that, diffusion models encode different parts of image information into different noise levels.
+At low noise level, the model tends to solve an image enhancement task, while at high noise level, the model is only suitable for extracting rough image morphology and layout [21].
+Intuitively, image understanding needs to take clean images as input to maintain complete visual information.
+However, due to the nature of diffusion modeling, this will only invoke the hidden representations that are related to the image enhancement task according to its low noise level, resulting in insignificant help for image understanding.
 Adding some noise to the input image would help invoke more semantic representations, but at a cost of image information loss.
@@
 (A): Most AR-based methods adopt discrete image tokens produced by VQ-VAEs.
-The VQ operation bottlenecks the representation’s information capacity, resulting in information loss.
-(B): Concurrent transfusion-like methods combining image DiT with text AR.
+The VQ operation bottlenecks per-tokens information capacity, resulting in information loss or huge token numbers.
+(B): Transfusion-like methods combining image DiT with text AR.
 $X^{(i)}$ denotes the full image at noise level $i$.
@@
 However, only $h^{(0)}$ is utilized in image to text tasks.
-(C): Our MMAR paradigm is VQ-free, thus reducing the information loss related to VQ operation.
-It auto-regressively models the continuous image tokens $x_{1},...,x_{n}$.
-The auto-regressive nature guarantees that all the image hidden representations, denoted by $h_{i},i\in[1,n]$, are simultaneously optimized for both $P(I)$ and $P(T|I)$, ensuring complete usage of image modeling capability for visual understanding tasks.
-In summary, how to take in and utilize the complete information of the continuous image modality is the major pain point of joint image-text probabilistic modeling.
+(C): Our MMAR paradigm utilizes continuous image tokens $x_{1},...,x_{n}$, enjoying efficient and lossless token compression.
+Its auto-regressive nature guarantees that all the image hidden representations, denoted by $h_{i},i\in[1,n]$, are simultaneously optimized for both $P(I)$ and $P(T|I)$, ensuring complete usage of image modeling capability for visual understanding tasks.
+In summary, how to efficiently take in and utilize the complete information of the continuous image modality is the major pain point of joint image-text probabilistic modeling.
 This is the challenge that our work is trying to address.
-The method proposed in this paper, MMAR, belongs to the image-text joint AR methodology, thus naturally avoids the challenge of fusing diffusion and AR, as mentioned previously.
-Different from other image-text joint AR methods, our method is built upon continuous image representation that can faithfully reconstruct the original image, rather than the discretized lossy image representation.
-To model the continuous probability distribution within the auto-regressive paradigm, we refer to MAR (Li et al., 2024b), introducing a simple mlp diffusion sampler to sample continuous image tokens given the transformer backbone output as its conditional information.
-This technique can also be regarded as disentangling the diffusion process from the auto-regressive backbone in the transfusion-like methods (Zhou et al., 2024; Zhao et al., 2024), which in turn focuses on learning complete visual representation that can guide the entire diffusion sampling process of image tokens.
-Different from MAR (Li et al., 2024b), we leverage LLMs to learn much more diverse distribution of the large-scale image-text data.
+As shown in Fig.
+2, the method proposed in this paper, MMAR, belongs to the image-text joint AR methodology, which ensures the complete use of image modeling capability by its auto-regressive nature.
+Different from other image-text joint AR methods, our method takes in continuous image tokens that can efficiently maintain the complete image information.
+To model the continuous probability distribution within the auto-regressive paradigm, we refer to MAR [30], introducing a simple mlp diffusion sampler to sample continuous image tokens given the transformer backbone output as its conditional information.
+This technique can also be regarded as disentangling the diffusion process from the auto-regressive backbone in the transfusion-like methods [71, 70], which in turn focuses on learning complete visual representation that can guide the entire diffusion sampling process of image tokens.
+Different from MAR [30], we leverage LLMs to learn much more diverse distribution of the large-scale image-text data.
 In practice, we find that low-precision training strategies for contemporary LLMs introduce non-negligible numerical error term in the diffusion sampler, which substantially impedes the convergence.
-We carefully analyze the numerical error term and find that the only way to minimize it is to properly parameterize the diffusion model.
+We carefully analyze the numerical error term and find that a effective way to minimize it is to properly parameterize the diffusion model.
 Therefore, we theoretically derive an optimal diffusion parameterization to solve this issue.
-Additionally, we find that visual understanding and generation prefer different experimental settings.
+Additionally, visual understanding and generation prefer different experimental settings.
 To balance these two tasks in a unified model, we propose a two-stage training method, which firstly improves visual understanding ability with large, mid quality data, then boost visual generation ability with high quality data.
@@
 Our contributions are the following three folds:
-  • We identified why previous joint probabilistic models for both image generation and understanding suffer image information loss with a series of experiments.
-  • We proposed a novel modeling framework that solves the issue by disentangling diffusion process from auto-regressive backbone, and achieved significant performance improvement for image understanding task.
+  • We identified why previous joint probabilistic models for both image generation and understanding suffer image information loss.
+  • We are the first to combine continuous image representation and discrete text representation into a unified auto-regressive probabilistic modeling framework.
+Our framework successfully overcomes information loss while keeping high efficiency, leading to significant performance gains on various image understanding tasks.
   • We proposed two training techniques crucial for training the model: one solves numerical error issue under low-precision training setting with theoretical guarantee, the other balances image generation and understanding tasks goals.
@@
 2.1 Multi-modal Large Language Models
-Since LLMs demonstrated open-domain conversational capabilities, more and more research has be focused on how to introduce visual information into large language models to achieve open-domain visual understanding.
-Pioneering works such as BLIP-2 (Li et al., 2023b), MiniGPT4 (Zhu et al., 2024) and LLaVA (Liu et al., 2023a) use trainable connector modules such as qformer or mlp to align image representations to the input space of LLM, making open-domain visual question answering possible.
-Recently, thanks to innovations in network architecture (Dong et al., 2024b; Tong et al., 2024), training strategy (Chen et al., 2024) and the support for dynamic resolution input (Hu et al., 2024; Li et al., 2024a), the visual understanding performance of large multi-modal models have been greatly improved.
+Since LLMs demonstrated open-domain conversational capabilities, more and more research has been focused on how to introduce visual information into large language models to achieve open-domain visual understanding.
+Pioneering works such as BLIP-2 [29], MiniGPT4 [72] and LLaVA [33] use trainable connector modules such as qformer or mlp to align image representations to the input space of LLM, making open-domain visual question answering possible.
+Recently, thanks to innovations in network architecture [10, 58], training strategy [8] and the support for dynamic resolution input [23, 27], the visual understanding performance of large multi-modal models have been greatly improved.
 These works focus on the alignment of image representations to text representations, only achieving $p(T|I)$ without incorporating the image’s own distribution $p(I)$ into the modeling capabilities of the model.
@@
 Auto-regressive methods represent a dominant paradigm in this domain, typically requiring discrete representations for both input and output.
-For images, this necessitates encoding them into discrete codes using a VQVAE (Esser et al., 2021; Kondratyuk et al., 2024).
-While recent works demonstrate that auto-regressive methods based on discrete image tokens can generate high-quality images (Sun et al., 2024a), the discretization of image representation acts as an information bottleneck, limiting the modeling accuracy of the image distribution.
-Recent efforts have shown that auto-regressive probabilistic modeling can be achieved without relying on discrete representations (Tschannen et al., 2023; Li et al., 2024b).
-For instance, MAR (Li et al., 2024b) replaces traditional logits with diffusion heads, enabling probabilistic sampling of continuous representations within an auto-regressive framework.
+For images, this necessitates encoding them into discrete codes using a VQVAE [13, 25].
+While recent works demonstrate that auto-regressive methods based on discrete image tokens can generate high-quality images [53], the discretization of image representation acts as an information bottleneck, limiting the modeling accuracy of the image distribution.
+Recent efforts have shown that auto-regressive probabilistic modeling can be achieved without relying on discrete representations [59, 30].
+For instance, MAR [30] replaces traditional logits with diffusion heads, enabling probabilistic sampling of continuous representations within an auto-regressive framework.
 This paper introduces continuous representation auto-regressive probability modeling to MLLMs, mitigating information loss caused by quantization and achieving theoretically lossless joint image-text probability modeling.
@@
 Early works in this area adopted a modular approach, bridging pre-trained models for visual understanding and visual generation through intermediate representations to achieve combined image understanding and generation.
-Notable examples include EMU (Sun et al., 2024c, b) and SEED-X (Ge et al., 2024).
+Notable examples include EMU [55, 54] and SEED-X [16].
 These works, however, are not considered probabilistic modeling because they aim at modeling the mean value of representations like CLIP or other intermediate representations rather than modeling the true image distribution.
 This limitation leads to the inadequate image space exploration, and thus hinders the attainment of high-quality generative and understanding performance.
-Another line of research adheres to the paradigm of probabilistic modeling (Team, 2024; He et al., 2024; Xie et al., 2024; Wu et al., 2024; Zhou et al., 2024; Zhao et al., 2024).
-These approaches can be categorized into three types based on whether the image representations are discrete and the modeling method of the image part : (i) Discrete auto-regressive Methods: Examples include Chameleon (Team, 2024), MARS (He et al., 2024), and VILA-U (Wu et al., 2024).
+Another line of research adheres to the paradigm of probabilistic modeling [57, 19, 62, 61, 71, 70].
+These approaches can be categorized into three types based on whether the image representations are discrete and the modeling method of the image part : (i) Discrete auto-regressive methods: Examples include Chameleon [57], MARS [19], VILA-U [61], and EMU-3 [60].
 These methods discretize image representations and then model images and text tokens using a unified auto-regressive transformer.
-(ii) Discrete Diffusion Methods: An example is Show-o (Xie et al., 2024).
+(ii) Discrete diffusion methods: An example is Show-o [62].
 These methods discretize image tokens and model them with text tokens using a unified transformer through a discrete diffusion approach.
-(iii) Continuous Diffusion Methods: Examples include Transfusion (Xie et al., 2024) and MonoFormer (Zhao et al., 2024).
+(iii) Continuous diffusion methods: Examples include Transfusion [71] and MonoFormer [70].
 These methods do not discretize image representations but directly employ continuous diffusion transformers to model image tokens along with text tokens using a unified transformer.
@@
 It belongs to the continuous auto-regressive method category, which does not require discretizing image representations.
-Instead, it predicts the continuous distribution of image tokens using an auto-regressive approach and models them alongside text tokens within a unified transformer, as shown in fig.1.
+Instead, it predicts the continuous distribution of image tokens using an auto-regressive approach and models them alongside text tokens within a unified transformer.
 3 Method
@@
 In the realm of natural language processing (NLP), the sequence $\mathbf{x}$ is solely made of discrete text tokens.
-As a result, most modern large language models (LLMs) parameterize $p_{\theta}(x_{i}|x_{<i})$ into a categorical distribution, which can be explicitly represented by the softmax activation on a set of logits predicted by a decoder-only transformer (Radford et al., 2019; Brown et al., 2020) $f_{\theta}(\cdot)$ followed by a linear LM head $H_{\theta}(\cdot)$:
+As a result, most modern large language models (LLMs) parameterize $p_{\theta}(x_{i}|x_{<i})$ into a categorical distribution, which can be explicitly represented by the softmax activation on a set of logits predicted by a decoder-only transformer [44, 1] $f_{\theta}(\cdot)$ followed by a linear LM head $H_{\theta}(\cdot)$:
 $p_{\theta}(x_{i}|x_{<i})=\mathrm{softmax}(H_{\theta}(f_{\theta}(x_{<i}))).$ (2)
@@
 Therefore, a protocol for parameterizing $p_{\theta}(x_{i}|x_{<i})$ of the continuous image tokens is required.
-Inspired by MAR (Li et al., 2024b), we train a diffusion model to achieve this.
+Inspired by MAR [30], we train a diffusion model to achieve this.
 The diffusion model takes vector $z_{i}=f_{\theta}(x_{<i})$ as the conditional input, and samples $x_{i}$ by gradually denoising from a randomly sampled gaussian noise.
 To optimize the diffusion model for continuous image token sampling, a diffusion loss can be applied, which acts as the upper-bound of the negative log likelihood.
-A typical diffusion loss can be written as follows, which is seen in MAR (Li et al., 2024b):
-$L(x_{i})=\mathbb{E}_{x_{i},\epsilon,t}[w_{t}\cdot||\mathbf{\epsilon}-\mathbf{\epsilon}_{\theta}(\sqrt{\bar{\alpha}_{t}}x_{i}+\sqrt{1-\bar{\alpha}_{t}}\mathbf{\epsilon},t,z_{i})||^{2}]\geq-\log p_{\theta}(x_{i}|x_{<i})+C,$ (3)
+A typical diffusion loss can be written as follows, which is seen in MAR [30]:
+$L_{i}$ $=\mathbb{E}_{x_{i},\epsilon,t}[w_{t}\cdot||\mathbf{\epsilon}-\mathbf{\epsilon}_{\theta}(\sqrt{\bar{\alpha}_{t}}x_{i}+\sqrt{1-\bar{\alpha}_{t}}\mathbf{\epsilon},t,z_{i})||^{2}]$
+$\geq-\log p_{\theta}(x_{i}|x_{<i})+C,$ (3)
 where $w_{t}$ is the loss weight that balances the loss for different timesteps, and $\bar{\alpha}_{t}$ indicates the noise schedule of the forward diffusion process.
-In this way, minimizing the diffusion loss is equivalent to maximizing the log likelihood of image data.
-The overall loss for joint image-text probabilistic modeling can be written as follows:
-$L=\sum_{i\in I_{img}}L(x_{i})-\sum_{i\in I_{txt}}\log p_{\theta}(x_{i}|x_{<i}),$ (4)
+In this way, minimizing the diffusion loss is equivalent to maximizing the log likelihood of image data conditioned on its preceding text guidance.
+The overall loss for joint image-text probabilistic modeling can be written as follows, which is an upper-bound of the negative log likelihood of the multi-modal data $\mathbf{x}$:
+$L=\sum_{i\in I_{img}}L_{i}-\sum_{i\in I_{txt}}\log p_{\theta}(x_{i}|x_{<i})\geq-\log p_{\theta}(\mathbf{x})+C,$ (4)
 where $I_{img}$ and $I_{txt}$ indicate the indices of image tokens and text tokens, respectively.
@@
 In the era of large language models, training with low precision data type, such as bfloat16, has become increasingly popular.
-However, the training and inference process of a diffusion model is relatively sensitive to numerical precision.
+However, the training and inference process of a diffusion model is relatively sensitive to numerical precision, we provide direct empirical demonstration in Appendix 6.8.
 Moreover, in an auto-regressive framework, the image tokens are sampled sequentially, requiring even more precise sampling for each image token to reduce the overall error accumulation.
 Therefore, handling the numerical error in the diffusion process modeling should be emphasized when integrating diffusion loss into LLMs.
+Figure 3: The overview of MMAR with two stage image expert training strategy.
 From the example below, we can clearly illustrate the effect of the numerical error in diffusion models.
-In eq.3, the diffusion model is parameterized as $\mathbf{\epsilon}_{\theta}(\sqrt{\bar{\alpha}_{t}}x_{i}+\sqrt{1-\bar{\alpha}_{t}}\mathbf{\epsilon},t,z_{i})$, known as “$\epsilon$-prediction”, predicting the noise $\epsilon$ that is added to the data.
+In Eq.
+3, the diffusion model is parameterized as $\mathbf{\epsilon}_{\theta}(\sqrt{\bar{\alpha}_{t}}x_{i}+\sqrt{1-\bar{\alpha}_{t}}\mathbf{\epsilon},t,z_{i})$, known as “$\epsilon$-prediction”, predicting the noise $\epsilon$ that is added to the data.
 Floating-point representation causes numerical error’s amplitude proportional to a number’s magnitude.
 This can be modeled by scaling the prediction by a factor of $(1+\delta)$, where $\delta$ is the relative error.
-In this way, we can write DDIM (Song et al., 2021) sampling with numerical error as follows:
-$\tilde{x}^{(t-1)}=$ $\sqrt{\bar{\alpha}_{t-1}}\left(\frac{x^{(t)}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(x^{(t)},t,z_{i})(1+\delta)}{\sqrt{\bar{\alpha}_{t}}}\right)+\sqrt{1-\bar{\alpha}_{t-1}}\epsilon_{\theta}(x^{(t)},t,z_{i})(1+\delta).$ (5)
+In this way, we can write DDIM [51] sampling with numerical error as follows:
+$\tilde{x}^{(t-1)}=$ $\sqrt{\bar{\alpha}_{t-1}}\left(\frac{x^{(t)}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(x^{(t)},t,z_{i})(1+\delta)}{\sqrt{\bar{\alpha}_{t}}}\right)$
+$+\sqrt{1-\bar{\alpha}_{t-1}}\epsilon_{\theta}(x^{(t)},t,z_{i})(1+\delta).$ (5)
 Further separating the numerical error term from the ideal DDIM sampling process, we get:
-$\tilde{x}^{(t-1)}=x^{(t-1)}+(\sqrt{1-\bar{\alpha}_{t-1}}-\frac{\sqrt{\bar{\alpha}_{t-1}}}{\sqrt{\bar{\alpha}_{t}}}\sqrt{1-\bar{\alpha}_{t}})\epsilon_{\theta}(x^{(t)},t,z_{i})\delta,$ (6)
-where the first term $x^{(t-1)}$ is the ideal DDIM sampling term, and the second term is the numerical error term.
-Assuming $\epsilon_{\theta}(x^{(t)},t,z_{i})$ follows a standard normal distribution, the standard deviation of numerical error term can be calculated as $|\sqrt{1-\bar{\alpha}_{t-1}}-\frac{\sqrt{\bar{\alpha}_{t-1}}}{\sqrt{\bar{\alpha}_{t}}}\sqrt{1-\bar{\alpha}_{t}}|\delta$.
+$\tilde{x}^{(t-1)}=$ $(\sqrt{1-\bar{\alpha}_{t-1}}-\frac{\sqrt{\bar{\alpha}_{t-1}}}{\sqrt{\bar{\alpha}_{t}}}\sqrt{1-\bar{\alpha}_{t}})\epsilon_{\theta}(x^{(t)},t,z_{i})\delta$
+$+x^{(t-1)},$ (6)
+where the first term is the numerical error term, and the second term $x^{(t-1)}$ is the ideal DDIM sampling term.
+Let $\sigma_{t}$ be the standard deviation of $\epsilon_{\theta}(x^{(t)},t,z_{i})$, then the standard deviation of numerical error term can be calculated as $|\sqrt{1-\bar{\alpha}_{t-1}}-\frac{\sqrt{\bar{\alpha}_{t-1}}}{\sqrt{\bar{\alpha}_{t}}}\sqrt{1-\bar{\alpha}_{t}}|\delta\sigma_{t}$.
 When the signal-to-noise ratio (SNR) is high, i.e.
@@
 Our goal is to minimize the numerical error term.
-To achieve this, we analyze all the factors that can determine the numerical error in Appendix A.2, and find that the only solution is to parameterize the diffusion model properly.
-By solving the numerical error minimization problem, we conclude that the v-prediction parameterization (Salimans & Ho, 2022) is the desired optimal parameterization method.
+To achieve this, we analyze most of the factors that can determine the numerical error in Appendix 6.2, and find that an effective solution is to parameterize the diffusion model properly.
+By solving the numerical error minimization problem, we conclude that the v-prediction parameterization [49] is the desired optimal parameterization method.
 Note that v-prediction is initially proposed for the efficient distillation of diffusion models, rather than reducing the numerical error of diffusion models.
 To the best of our knowledge, our work is the first to derive v-prediction parameterization from the first principle of minimizing the numerical error in diffusion models.
-For more details, see Appendix A.2.
+For more details, see Appendix 6.2.
 Under v-prediction parameterization, the model predicts a linear combination of data $x_{i}$ and noise $\epsilon$:
-$v^{(t)}_{i}=\sqrt{\bar{\alpha}_{t}}\mathbf{\epsilon}-\sqrt{1-\bar{\alpha}_{t}}x_{i}$ (7)
+$v^{(t)}_{i}=\sqrt{\bar{\alpha}_{t}}\mathbf{\epsilon}-\sqrt{1-\bar{\alpha}_{t}}x_{i}.$ (7)
 We therefore re-write Eq.3 into the v-prediction form, and set the loss weight $w_{t}$ to 1 for simplicity:
-$L(x_{i})=\mathbb{E}_{x_{i},\epsilon,t}[||v_{i}^{(t)}-v_{\theta}(\sqrt{\bar{\alpha}_{t}}x_{i}+\sqrt{1-\bar{\alpha}_{t}}\mathbf{\epsilon},t,z_{i})||^{2}].$ (8)
-3.3 Model Architecture and Training Strategy
-Figure 2: The overview of MMAR with two stage image expert training strategy.
-3.3.1 Pipeline
-To jointly model the probabilities of images and text at the token level, we employ an image tokenizer to extract image tokens and randomly mask some of them for image generation training, as depicted in fig.
-2.
-Given the substantial distribution gap between images and text, directly scaling image tokens to match the LLM’s dimensions would pose a challenge to joint image-text probability modeling.
-To address this, we introduce a Latent Encoding process utilizing an EmbeddingViT module for the unmasked image tokens, which simplifies the modeling task.
-To facilitate the auto-regressive training of image tokens, we concatenate the known image tokens with the masked tokens and append the position embedding corresponding to their original positions.
-These tokens, after being processed by the VisProjector, are concatenated with the text tokens to generate the final image-text sequence.
-Then, an LLM is used to process the image-text sequence, whose output $z$ integrates information from both the preceding text and the unmasked image tokens.
-For image output, $z$ acts as a conditional input for the Diffusion MLP, which predicts the $v$ values of the noisy masked image tokens.
-The diffusion loss, as depicted in eq.8, is applied, enabling us to model the probability of the continuous image tokens.
-For text output, a linear lm-head processes $z$ to predict the next token’s logits, and the corresponding cross entropy loss is applied, allowing us to model the probability of the text tokens.
-During image generation, to ensure consistency with the training scenario, we initialize all image tokens as mask tokens and randomly generate a set of image position sequences.
-Following a left-to-right order, we extract N condition tokens $z$ at each step and input them into the Diffusion MLP to generate the corresponding image tokens.
-Next, we assign these image tokens back to their respective positions in the original image token sequence, iterating until the entire image token set is obtained.
-Ultimately, we decode the image tokens into images using the image tokenizer, yielding the generated images.
-3.3.2 Model Architecture
-Tokenizer
-MMAR employs the publicly available KL-16 image tokenizer from LDM (Rombach et al., 2022).
-This tokenizer processes an 256 $\times$ 256 image into 256 image tokens, each with 16 channels, and it is kept frozen during training.
-EmbeddingViT
-To achieve a more profound encoding of image tokens, we introduce the EmbeddingViT module, a ViT(Dosovitskiy et al., 2021) encoder with 16 layers and 1024 hidden state channels, processing image tokens into visual embeddings with stronger context awareness.
-Moreover, as the input consists of randomly shuffled image tokens, we integrate a learnable position embedding for each image token in EmbeddingViT, corresponding to its original position.
-LLM
-We initialize our LLM using parameters from the open-source QWen2 series models (Yang et al., 2024).
-To preserve text modeling capabilities, we keep the LLM parameters fixed during training and add PLoRA (Dong et al., 2024b) as the image expert, where only the image tokens in the input pass through the introduced LoRA (Hu et al., 2022) adapters.
-The PLoRA is applied for each linear layer within the original LLM.
-Considering the time cost, our ablation study employs QWen2-0.5B-Instruct, with $r=512$ for PLoRA.
-Furthermore, we use QWen2-7B-Instruct to explore our method’s scale-up capability with $r=1280$ for PLoRA.
-The corresponding MMAR models are denoted by MMAR-0.5B and MMAR-7B, respectively.
-In the LLM, we adopt a bidirectional attention mechanism to enhance interaction among image tokens, rendering all image tokens mutually visible, as depicted in fig.
-2.
-Meanwhile, the text portion retains causal attention.
-Additionally, to prevent interference with the auto-regressive training of image tokens in random order, we assign the same ROPE position id to all image tokens within all LLM layers.
-This strategy not only ensures the random order regression of image tokens but also mitigates the issue of text concentrating on closer image tokens in long text scenarios.
-Diffusion MLP
-Inspired by MAR, we employ a simple MLP to predict $v^{(t)}$.
-This MLP consists of a series of residual blocks, each comprising AdaLN (Peebles & Xie, 2023), a linear layer, SiLU, and another linear layer.
-The condition vector $z$ is added to the diffusion time embedding and incorporated through AdaLN.
-MMAR-0.5B utilizes a Diffusion MLP with 8 residual blocks and 1024 channels, while MMAR-7B employs a larger variant with 12 residual blocks and 2048 channels.
-3.3.3 Training Strategy
-Multi-Task Training
-To accomplish joint image-text modeling, we simultaneously conducted text-to-image, image-to-text, and unconditional image generation tasks during training.
-In the image-to-text task, no mask tokens are assigned to the image part, allowing us to model $P(T|I)$ using the complete image tokens.
-Furthermore, the allocation ratio of text-to-image and unconditional image generation tasks is set to 9:1, facilitating efficient use of the cfg technique at the inference stage.
-To maintain a balance between tasks, we intuitively set the sample allocation ratio of image generation tasks and image understanding tasks to 1:1.
-Two Stage Training
+$L_{i}=\mathbb{E}_{x_{i},\epsilon,t}[||v_{i}^{(t)}-v_{\theta}(\sqrt{\bar{\alpha}_{t}}x_{i}+\sqrt{1-\bar{\alpha}_{t}}\mathbf{\epsilon},t,z_{i})||^{2}].$ (8)
+3.3 Implementation
+3.3.1 Model Design
+Overall, our model follows the auto-regressive (AR) modeling paradigm for both image and text.
+For the text part, we leverage a pre-trained decoder-only LLM [63] for the modeling of text probability $P(T)$.
+While for the image part, we pick the masked AR implementation [30], which has been demonstrated to perform much better than decoder-only transformer implementation, meanwhile maintaining the form of $p_{\theta}(x_{i}|x_{<i})$ modeling.
+Our design goal is to embed the masked AR into the pre-trained LLM to realize a unified $f_{\theta}(\cdot)$ for multi-modal input and output.
+Generally, a masked AR implementation includes a bi-directional encoder, which encodes the known image tokens $x_{<i}$, and a bi-directional decoder, which decodes the masked image token $x_{i}$ given $x_{<i}$ and the mask position embeddings.
+Our approach is shown in Fig.
+3.
+Firstly, we introduce an EmbeddingViT module to serve as the role of the encoder.
+Secondly, we directly embed the decoder into the LLM.
+This is achieved by adding PLoRA-based [10] image expert into the LLM, and setting the attention of the image region to be bi-directional.
+We also set the same ROPE position ID for all image tokens to ensure that only the mask position embedding decides the image token position information.
+In this way, the decoder-only LLM behaves just like the masked AR decoder when the PLoRA is activated.
+Meanwhile, thanks to the pre-trained LLM, the decoding of image tokens can be conditioned on the preceding text.
+To calculate the image diffusion loss, a Diffusion MLP module is adopted.
+This module takes the embedding of the image part output by the LLM, denoted by $z_{i}=f_{\theta}(x_{<i})$, as a conditional input to predict $v_{i}^{(t)}$ according to Eq.
+7, ultimately achieving the loss of Eq.
+8.
+For detailed module specifics, please refer to Appendix 6.1.
+During training, we simultaneously learn two tasks: image understanding and image generation.
+For image understanding task, the sequence is arranged as [img][text], and we perform text auto-regressive training under the condition of all image tokens, achieving $P(T|I)$.
+For conditional image generation task, the sequence is arranged as [text][img], and we perform image masked AR training under the condition of preceding text tokens, achieving $P(I|T)$.
+Specifically, the image tokens are randomly masked with a controlled mask ratio decided by the training strategy.
+For unconditional image generation task, the sequence is arranged as [img] only, and we perform image masked AR training without any preceding text conditions, achieving $P(I)$.
+This also allows the application of Classifier-Free Guidance (CFG) techniques.
+Combining $P(I|T)$, $P(T|I)$, $P(I)$, and $P(T)$, we achieve the modeling of $P(I,T)=P(I|T)P(T)=P(T|I)P(I)$, form both text to image and image to text direction.
+3.3.2 Two Stage Training
+Image understanding tasks are often applied in various complex scenarios.
+Therefore, the training requires highly diverse data, including but not limited to blurred images, small-sized images, images with greatly differing aspect ratios, etc.
+However, image generation tasks, on the other hand, place higher demands on data clarity, resolution, and even aesthetic criteria is one of the ways to evaluate generative performance.
 To achieve a balance between image generation and understanding capabilities, our training process is carried out in two stages.
-The first stage utilizes large-scale, mid-quality data (as illustrated in fig.
-2) to enhance the diversity of the model’s data distribution (Hu et al., 2024).
+The first stage, called Image Expert Pretraining stage, utilizes large-scale, mid-quality data (as illustrated in Fig.
+3) to enhance the diversity of the model’s data distribution [23].
 By initially modeling this diverse data, we strengthen the model’s understanding capabilities.
-In the second stage, we employ a smaller volume of high-quality data to further improve the image generation capacity and refine the model’s comprehension of images.
-We observe that when the training dataset is excessively large, preventing the model from iterating over each image hundreds of times, maintaining an image mask ratio within the range of [0.7, 1], as done in MAR, hinders the model’s ability to generate coherent image tokens.
-Consequently, in the second stage, we adjust the mask ratio to (0, 1].
-The example is showed in Appendix A.5.
+In the second stage, called Image Expert Fine-tuning stage, we employ a smaller volume of high-quality data to further improve the image generation capacity and refine the model’s comprehension of images.
+Surprisingly, we find that the images generated by the model at the end of the first training stage all have holes, as shown in Fig.
+4.
+We control the step-by-step generation of images tokens and ultimately find that the appearance of holes always occurs within the last thirty percent of the generation sequence.
+And this is because the first stage uses a [0.7, 1.0] mask ratio, which in the context of fewer training epochs, results in the model still not adapting to the image completion scene task of the last thirty percent.
+To address this issue, we set the lower bound of the mask ratio to 0 in the second stage and adjust the variance of the mask ratio distribution to enable the model to learn image generation tasks in various scenarios as efficiently as possible.
 The probability density curve of the two mask ratios are shown on the right side of Fig.
-2.
-Additionally, due to the small size of the diffusion mlp and the limited amount of high-quality data, we increased the number of timestep samples for the given $z$ to enhance learning efficiency.
-To further mitigate performance fluctuations in the model, we employed an exponential moving average (EMA) method with a momentum of 0.9999 in both stages.
-3.3.4 Inference Strategy
-Classifier-Free Guidance
-During the inference stage, the model performs both text-based and non-text-based image generation tasks.
-The provided conditions are represented as $z_{c}$ and $z_{u}$, and the predicted $v$ is given by: $v=v_{\theta}(x^{(t)}|t,z_{u})+\omega*(v_{\theta}(x^{(t)}|t,z_{c})-v_{\theta}(x^{(t)}|t,z_{u}))$ (which has the same effect as $\epsilon=\epsilon_{u}+\omega*(\epsilon_{c}-\epsilon_{u})$ , with the mathematical derivation process detailed in the Appendix A.4), where $\omega$ denotes the guidance scale.
-Our method has been experimentally validated to support a large guidance scale, as compared to the noise prediction training approach.
-We hypothesize that this is due to the strong conditioning provided by $z$, which, coupled with the auto-regressive (AR) process, necessitates reducing accumulated error.
-Thus, each token must be generated as accurately as possible, a requirement that can be fulfilled with a large guidance scale.
-Mixed-Precision Inference
-To enhance numerical stability during image generation, particularly when using small step sizes in the DDPM sampling process, we perform the model’s forward pass in bfloat16 (matching the training precision) but cast the output to float32 before DDPM sampling.
-This mitigates potential rounding error without a significant computational overhead, therefore improving the sampling accuracy efficiently.
+3.
+Ultimately, this achieves a rapid improvement in the model’s generative capabilities within a limited number of training epochs.
+Figure 4: Training stage 2 with (0,1] mask ratio range addresses the issue of generating holes of training stage 1.
+Table 1: Comparison on visual understanding benchmarks.
+MMAR surpasses other joint image-text probabilistic models even with a small resolution of 256x256, approaching the performance of traditional MLLMs like LLaVA, which employ pre-trained CLIP vision encoder.
+Type Method LLM V-Token Res. MMB MMEP POPE SEED MM-Vet AVE@18Und.
+Und. Only LLaVA-1.5 [34] Vicuna-1.5-7B CLIP 336 64.3 1510.7 85.9 58.6 31.1 47.08
+Und. & Gen. Non-prob. EMU-2 [55] LLaMA-13B CLIP 448 62.8 48.5
+SEED-X [16] LLaMA-13B CLIP dynamic 75.4 1435.7 84.2
+Joint Prob. Models Chameleon-7B [57] 7B from scratch vq-vae 512 13.32 125.8 30.86 34.61 7.34 18.34
+Transfusion* Qwen-2-0.5B vae 256 29.47 594.3 66.90 42.40 13.90 28.26
+Show-o [62] Phi-1.5B CLIP 336 42.44 1182.7 84.50 51.61 20.87 33.06
+VILA-U [61] LLaMA-2-7B vq-vae 256 1336.2 83.9 56.3 27.7
+EMU-3 [60] 8B from scratch vq-vae 512 58.5 85.2 68.2 37.2
+MMAR-0.5B Qwen-2-0.5B vae 256 48.45 882.1 70.74 55.70 18.49 34.56
+MMAR-7B Qwen-2-7B vae 256 70.45 1486.9 84.02 68.63 30.64 48.25
 4 Experiment
 4.1 Dataset
-We utilized the Capfusion-120M dataset for the image expert pretraining stage.
-This dataset is publicly accessible and comprises an extensive collection of web-based image-text pairs, designed to optimize noisy captions (Yu et al., 2023).
+We utilized the Capfusion-120M dataset [64] for the image expert pretraining stage.
+This dataset is publicly accessible and comprises an extensive collection of web-based image-text pairs, designed to optimize noisy captions.
 In an effort to further improve the quality of the content generated, we executed a random sampling of 20M data points from the Capfusion dataset during our image expert fine-tuning stage.
-This was supplemented with a high-quality mixed dataset that included ImageNet-1K-1.2M (Deng et al., 2009), CC12M (Changpinyo et al., 2021), and LAION-aesthetics-12M [^1] https://huggingface.co/datasets/dclure/laion-aesthetics-12m-umap.
-We employ the open-source InternVL2-8B for recaptioning the CC12M and laion-aesthetics-12m datasets in English.
-Following LLaVA-v1.5 (Liu et al., 2023a), we use LLaVA-v1.5-mix-665K data for instruction tuning before each performance test for visual understanding.
+This was supplemented with a high-quality mixed dataset that included CC12M [2], and LAION-aesthetics-12M [^1] https://huggingface.co/datasets/dclure/laion-aesthetics-12m-umap.
+We employ the open-source InternVL2-8B [7] for recaptioning the CC12M and laion-aesthetics-12m datasets in English.
+Following LLaVA-v1.5 [33], we use LLaVA-v1.5-mix-665K data for instruction tuning before each performance test for visual understanding.
 4.2 Implementation Details
-By default, we utilized the AdamW optimizer with betas (0.9, 0.95).
-The weight decay was consistently maintained in proportion to the learning rate.
-During the first training stage, a learning rate of 5e-5 was employed for a total of 4 epochs, with an initial warm-up phase of 0.5 epochs, followed by maintaining the learning rate at 5e-5.
-In the second stage, the 0.5B model maintained a 5e-5 learning rate and a 0.5 epoch warm-up, with training lasting for 3 epochs.
-For the larger 7B model, we initially applied a 2e-6 learning rate for 1.5 epochs, before transitioning to a 1e-6 learning rate for an additional 1.5 epochs.
-Furthermore, during the first stage, the total batch size for the smaller 0.5B model was 2496, while it was 1152 for the larger 7B model.
-In the second stage, the total batch size for the smaller 0.5B model was 768, and for the larger 7B model, it was 480.
-Notably, in both stages, only the PLoRA portion of the LLM parameters is trainable, and a consistent image resolution of 256x256 was used throughout.
-Our models utilize bfloat16 during training and inference.
-For DDPM sampling, we convert the model output to float32 for higher precision.
+By default, we used the AdamW [37] optimizer with betas (0.9, 0.95) and weight decay proportional to the learning rate.
+In the first training stage, a learning rate of 5e-5 was used for 4 epochs (with a 0.5-epoch warm-up) for the 0.5B model, and 2e-5 for 3 epochs (with a 0.5-epoch warm-up) for the 7B model.
+The second stage used 5e-5 for 3 epochs (with a 0.5-epoch warm-up) for the 0.5B model, and 2e-5 for 3 epochs (with a 0.5-epoch warm-up) for the 7B model.
+Batch sizes were 2496 for the 0.5B model and 1152 for the 7B model in the first stage, and 768 for the 0.5B model and 480 for the 7B model in the second stage.
+The image tokenizer and original pre-trained LLM parameters were frozen in both stages, while other parameters were trainable.
+A consistent image resolution of 256x256 and the LDM-KL-16 [47] tokenizer were used throughout.
 4.3 Comparison with Other Systems
-Table 1: Comparison on visual understanding benchmarks.
-MMAR surpasses other joint image-text probabilistic models by a large margin, even with a small resolution of 256x256, approaching the performance of traditional MLLMs like LLaVA, which employ pretrained CLIP vision encoder.
-Method LLM V-Token Res. MMB MMEP POPE SEED MM-Vet AVE@18Und.
-LLaVA-1.5 Vicuna-1.5-7B CLIP 336 64.3 1510.7 85.9 58.6 31.1 47.08
-EMU-2 LLaMA-13B CLIP 448 62.8 48.5
-SEED-X LLaMA-13B CLIP dynamic 75.4 1435.7 84.2
-DreamLLM Vicuna-7B CLIP 224 58.2 36.6
-Chameleon-7B 7B from scratch vq-vae 512 13.32 125.8 30.86 34.61 7.34 18.34
-Transfusion* Qwen-2-0.5B vae 256 29.47 594.3 66.90 42.40 13.90 28.26
-Show-o Phi-1.5B CLIP 336 42.44 1182.7 84.50 51.61 20.87 33.06
-VILA-U LLaMA-2-7B vq-vae 256 1336.2 83.9 56.3 27.7
-MMAR-0.5B Qwen-2-0.5B vae 256 48.45 882.1 70.74 55.70 18.49 34.56
-MMAR-7B Qwen-2-7B vae 256 66.32 1393.9 83.02 64.52 27.80 46.52
 Visual Understanding.
-As depicted in table 1, we thoroughly gauge MMAR’s performance in visual understanding, employing VLMEvalKit (Duan et al., 2024) to perform extensive evaluations on prevalent visual understanding benchmarks, encompassing a total of 18 such assessments (average score denoted by “AVE@18Und.” in table 1, for details see Appendix A.7.) including MMB (Liu et al., 2023b), MME (Fu et al., 2023), POPE (Li et al., 2023c), SEED (Li et al., 2023a), MM-Vet (Yu et al., 2024), among others.
-Our method outperforms other joint image-text probabilistic models by a large margin, including Chameleon-7B (Team, 2024), Show-o (Xie et al., 2024), VILA-U (Wu et al., 2024) and our re-implemented version of Transfusion (denoted by “Transfusion*”), approaching the performance of traditional MLLMs like LLaVA, which employ pretrained CLIP vision encoder.
-Even without using any pre-trained CLIP or diffusion models and with small resolution of 256 $\times$ 256, MMAR-7B presents comparable or even better performance when compared to methods using pre-trained clip and diffusion models, including EMU-2 (Sun et al., 2024c), SEED-X (Ge et al., 2024), and DreamLLM (Dong et al., 2024a).
+In Table 1, we employ VLMEvalKit [12] to perform extensive evaluations on prevalent visual understanding benchmarks, encompassing a total of 18 such assessments (average score denoted by “AVE@18Und.” in Table 1, for details see Appendix 6.6.) including MMB [35], MME [15], POPE [31], SEED [26], MM-Vet [65], among others.
+Our method outperforms other joint image-text probabilistic models by a large margin, including Chameleon-7B, Show-o, VILA-U and our re-implemented version of Transfusion (denoted by “Transfusion*”).
+With merely 256 image token input, it achieves comparable or better performance compared to EMU3, which needs 4096 tokens per image.
+Even without using any pre-trained CLIP or diffusion models and with small resolution of 256 $\times$ 256, MMAR-7B presents comparable or better performance when compared to methods rely on pre-trained clip and diffusion models, including understanding-only methods like LLaVA-1.5, and non-probabilistic joint image understanding and generation models like EMU-2 and SEED-X.
 Visual Generation.
-We showcase the zero-shot FID of MMAR, evaluated on the MSCOCO 30k dataset, in Table 2.
-Our model’s performance is discernibly on par with current robust generative models, a noteworthy achievement for MMAR given its minimal training epochs.
-Repeated exposure to identical data can substantially enhance a model’s generative quality, as exemplified by MAR, which trains for 400 epochs on the 1.2M ImageNet dataset, and Show-o, which undergoes approximately 49 epochs of training on the 35M dataset.
+We showcase the zero-shot FID [20] of the MMAR at the end of the second stage training, evaluated on the MSCOCO 30k dataset [32], in Table 3.
+It is worth noting that, even with only 3 epochs of training and using a less pure high-quality dataset due to the need to balance understanding performance, MMAR has been able to generate reasonable images, although it has not achieved very good generation effects.
+Furthermore, in order to verify that MMAR has the ability to generate high-quality images, we additionally train for 6 epochs on a pure high-quality image dataset as the third stage, where we replace Capfusion with JourneyDB [52] based on the second stage dataset.
+In the third stage, we only use image diffusion loss to further explore the model’s generative capabilities.
+We evaluate MMAR on FID@MJHQ-30K [28] and GenEval [17] benchmarks at the end of the third stage training, and show results in Table 4.
+It can be seen that after training with pure high-quality images, our model’s performance is discernibly on par with existing methods.
+Meanwhile, our image loss is still decreasing, which indicates that the image generation capability can still be improved.
+Notably, we also evaluated the visual understanding ability of the MMAR at the end of the third stage, and the overall average metric only decreased by 1%.
+For generated example images and visual understanding performance, please refer to Appendix 6.7 and 6.6.
+Table 2: Ablation study on MMAR
+[…]
+MMAR-0.5B 48.45 882.1 34.56 36.6
+w/ $\epsilon$-pred. 45.53 880.7 32.21 61.53
+show-O like (w/VQ) 37.54 618.2 29.70 66.26
+transfusion-like 29.47 594.3 28.26 95.38
+4.4 Ablation Study
+Table 2 demonstrates that switching to the more common $\epsilon$-prediction leads to a significant decrease in both visual understanding and generation quality, confirming the effectiveness of our optimal diffusion parameterization.
+To evaluate the efficacy of various image-text joint modeling techniques, we devise two distinct versions based on the MMAR framework: one employing LDM-VQ-16 tokenizer [47] for discrete token modeling (refer to show-O), and the other utilizing Transformer for diffusion modeling (refer to Transfusion).
+Implementation details are provided in the Appendix 6.1.
+Our test results are presented in Table 2.
+The transfusion-like version considerably underperforms the other two approaches in both understanding and generation aspects.
+This is attributed to the substantial loss of image information when jointly modeling image-text and operating with limited training epochs.
+In contrast, our full method consistently delivers superior results.
 Table 3: Comparison on MSCOCO Dataset
 […]
-Gen. Only DALL-E 12B 250M 27.50
-LDM 1.4B 400M 12.64
-DALL-E2 6.5B 650M 10.39
-Imagegen 3B 5000M+ 6.61
-Und. and Gen. w/ pre-trained Diff. CoDI - 400M 11.26
-SEED-X 17B - 12.68
-DreamLLM 7B - 8.76
-Joint Prob. Models Show-o 1.3B 35M 9.24
-Chanmeleon 7B - 29.6
-Transfusion  [^2] Both the Transfusion and Chanmeleon results are referenced from Table 3 in the paper ‘Transfusion: Predict the Next Token and Diffuse Images with One Multi-Modal Model.’ 7B - 16.8
+Gen. Only DALL-E [45] 12B 250M 27.50
+LDM [47] 1.4B 400M 12.64
+DALL-E2 [46] 6.5B 650M 10.39
+Imagegen [48] 3B 5000M+ 6.61
+Und. and Gen. w/ pre-trained Diff. CoDI [56] - 400M 11.26
+SEED-X [16] 17B - 12.68
+DreamLLM [9] 7B - 8.76
+Joint Prob. Models Show-o [62] 1.3B 35M 9.24
+Chanmeleon [57] 7B - 29.6
+Transfusion [71]  [^2] Both the Transfusion and Chanmeleon results are referenced from Table 3 in the paper ‘Transfusion: Predict the Next Token and Diffuse Images with One Multi-Modal Model.’ 7B - 16.8
 MMAR-0.5B 0.5B 145.2M 36.6
-MMAR-7B 7B 145.2M 17.1
-However, to bolster the model’s comprehension capabilities, exposure to diverse data is crucial rather than simply reiterating the same data.
-Consequently, we restrict our training to just 3 epochs for the second stage high-quality dataset.
-Despite these constraints, our model maintains competitive performance in generation, further substantiating the efficacy of our image-text joint modeling approach.
-The example images generated by our model are shown in the Appendix A.6.
-Scaling up with Model Size. From Table 1 and 2, we see that MMAR can scale up from 0.5B to at least 7B model size, with significant improvement of visual understanding and generation capability.
-4.4 Ablation Study
-We begin by evaluating the impact of our chosen diffusion parameterization method.
-Table 3 demonstrates that switching to the more common $\epsilon$-prediction leads to a significant decrease in both visual understanding and generation quality, confirming the effectiveness of our optimal diffusion parameterization.
-To evaluate the efficacy of various image-text joint modeling techniques, we devise two distinct versions based on the MMAR framework: one employing LDM-VQ-16 tokenizer (Rombach et al., 2022) for discrete token modeling, and the other utilizing Transformer for diffusion modeling (refer to Transfusion).
-Comprehensive implementation details are provided in the Appendix A.1.
-Our test results are presented in Table 3.
-The Transformer-based diffusion modeling version considerably underperforms the other two approaches in both understanding and generation aspects.
-This is attributed to the substantial loss of image information when jointly modeling image-text and operating with limited training epochs.
-In contrast, our method consistently delivers superior results.
-Table 3: Ablation study on MMAR
-[…]
-MMAR-0.5B(Full method) 48.45 882.1 70.74 55.70 18.49 34.56 36.6
-w/ $\epsilon$-pred. 45.53 880.7 71.14 53.72 17.98 32.21 61.53
-w/ VQ 37.54 618.2 66.98 44.93 14.45 29.70 66.26
-transfusion-like 29.47 594.3 66.90 42.40 13.90 28.26 95.38
+MMAR-7B 7B 145.2M 22.9
 4.5 Analysis
@@
 5 (A).
-Furthermore, to more effectively discern the loss discrepancies between the two techniques, we subtract the v-prediction curve from the $\epsilon$-prediction curve, yielding the yellow curve.
-For reference, we plot the theoretically predicted numerical error of the $\epsilon$-prediction model (see Appendix A.3) as the red curve.
+Furthermore, to effectively discern the loss discrepancies between the two techniques, we plot the difference between v-prediction and the $\epsilon$-prediction curve, yielding the yellow curve.
+For reference, we plot the theoretical numerical error of the $\epsilon$-prediction model (details in Appendix 6.3) as the red curve.
 The graph reveals that the loss of $\epsilon$-prediction model is consistently higher than v-prediction model.
@@
 The gap between yellow and red curve indicates that apart from its direct effect, numerical error also introduces optimization difficulty, hindering the loss convergence.
-Impact of CFG Scaling
-We select models from the second and fourth epochs of the first stage as starting points for the second stage, train them for 3 epochs, and then test the MSCOCO FID-30K under varying CFG intensities.
-As shown in fig.
-3 (B), our method achieves better FID scores as the CFG scale increases from 1 to 10.
-It is worth noting that most probabilistic generative models typically have a CFG scale between 1.5 and 5.
-Additionally, it is observed that a longer training duration in the first stage consistently results in better generation outcomes at all CFG scales.
-Scaling up with More Training Data
-As illustrated in fig.
-3 (C), we evaluate the average performance of MMAR-0.5B on 18 visual understanding benchmarks, following SFT training applied to the checkpoints generated during its training process.
-The blue background in the figure denotes the Image Expert Pretraining stage, while the green background signifies the Image Expert Fine-tuning stage.
-The curve reveals that, with an increasing number of training steps, i.e.
-more training data, the visual understanding performance of MMAR-0.5B consistently improves without reaching saturation.
-This finding highlights the exceptional scale-up capability of our MMAR model.
+Table 4: Comparison on MJHQ-30K Dataset and GenEval.
 […]
-Figure 3: The impact of v-prediction, CFG scaling, and training steps.
-4.6 Limitation
-Our method still requires further optimization in terms of image generation speed.
-Although the diffusion mlp is not involved in image understanding tasks, when applied to image generation tasks, we are compelled to use a 256-step token generation followed by a 100-step diffusion denoising process to ensure the quality of the generated images.
-This results in a generation time of nearly three minutes for a single 256x256 image using MMAR-7B.
-While it is possible to generate multiple images simultaneously by increasing the batch size, this does not fundamentally resolve the issue of prolonged generation time.
-We plan to address this in future work.
-5 Conclusion
-This paper proposes MMAR, a novel multimodal auto-regressive probabilistic modeling framework based on continuous image representations.
-It employs a standalone diffusion mlp at the image token level on top of pre-trained LLMs to facilitate theoretical lossless joint image-text probabilistic modeling.
-In practice, the low precision training of LLMs poses an non-negligible numerical error term to diffusion loss, causing optimization difficulty.
-This was addressed by deriving an optimal diffusion model parameterization.
-To balance the understanding and generation ability, a two-stage training strategy is introduced.
-During inference time, MMAR can tolerant extremely large CFG scale to generate high quality images.
-MMAR significantly demonstrates scaling-up laws with more data and larger model size.
-Extensive evaluations are conducted on 18 image understanding benchmarks, revealing that MMAR is the first joint image-text modeling framework that approaches comparable performance with traditional MLLMs that employ pretrained CLIP vision encoder, marking a significant step toward lossless joint probabilistic modeling of images and text.
-Appendix A Appendix
-A.1 Additional Implementation Details
-VQ
-Based on the MMAR-0.5B framework, we replace the Image Tokenizer from KL-16 to VQ-16.
-The image codes extracted using VQ-16 are then passed through a projector to increase the channel size to match the LLM’s hidden size.
+Gen. Only PixArt [3] 6.14 0.48
+SDXL [43] 9.55 0.55
+SD3 [14] - 0.68
+Und. and Gen. w/ pre-trained Diff. CoDI [56] - 0.31
+SEED-X [16] - 0.55
+Joint Prob. Models Show-o [62] 15.18 0.53
+VILA-U [61] 12.81 -
+Transfusion [71] - 0.63
+EMU3 [60] - 0.66
+MMAR-7B 15.6 0.51
+Scaling Law
+We conduct scaling up experiments by varying model size (N) and training token count (D), approximating FLOPs as 6ND.
+Fig.
+5 (B), (C) displays the performance on image understanding and generation benchmarks.
+The scaling laws are evident in both tasks, as shown by the dashed lines.
+[…]
+Figure 5: The impact of v-prediction and the scaling laws on the tasks of image understanding and generation.
+5 Discussion and Conclusion
+MMAR is the first to achieve the unified autoregressive probabilistic modeling for continuous image and discrete text representation.
+It addresses numerical instability and balances image comprehension and generation, thereby ensuring practicality and scalability.
+While masked autoregressive modeling serves as one of the design options for MMAR, our framework can be adaptable to future breakthroughs in continuous image autoregressive modeling.
+MMAR uses continuous image tokens, boosting efficiency sixteenfold over image discretization, matching or exceeding EMU-3’s image comprehension capabilities with fewer image tokens (4096 tokens/image to 256 tokens/image).
+This advantage grows with advanced continuous image tokenizers.
+And it would definitely benefit more when it comes to modeling of interleaved multi-image data and long videos.
+MMAR is currently in early exploration, being experimented on relatively limited resolution, data scale, model size, and did not utilize more advanced continuous image tokenizers.
+Despite this, MMAR shows promise through its scalable performance and strong comprehension/generation abilities.
+Future works include unlocking these limitations and developing more advanced continuous image AR implementation other than MAR.
+Acknowledgments
+This work is supported by the National Natural Science Foundation of China (NSFC) under Grants 62225207, 62436008 and 62306295. Additionally, we thank Ziheng Zhou for his helpful discussion.
+Supplementary Material
+6 Appendix
+6.1 Additional Implementation Details
+Table 5: Hyper-parameter settings for MMAR models
+MMAR-0.5B MMAR-7B
+Module Param. settings settings
+$N_{diff}$ 8 12
+$d_{mlp}$ 1024 2048
+Diffusion MLP $r_{mlp}$ 4 4
+Res. 256 256
+Image Tonkenizer $d$ 16 16
+$N_{ViT}$ 16 16
+EmbeddingViT $d_{ViT}$ 1024 1024
+$r$ 512 1280
+LLM’s PLoRA $\alpha$ 128 128
+Diffusion MLP.
+Inspired by MAR, we employ a simple MLP architecture to predict $v^{(t)}$, whose detailed architecture is shown in Fig.
+6.
+It consists of a main network and its input/output linear projection layers.
+The input projection converts the $d$-dimensional noisy latent $x^{(t)}$ into the mlp hidden dimension $d_{mlp}$, while the output projection converts the $d_{mlp}$-dimensional main network output back to $d$ dimensions.
+The main network is a stack of $N_{diff}$ residual blocks, each comprising an AdaLN [42], followed by a two-layer MLP activated by SiLU.
+The expand ratio of each MLP is denoted by $r_{mlp}$, which means its first linear layer projects from $d_{mlp}$ to the dimension of $d_{mlp}\times r_{mlp}$.
+The condition vector $z$ is added to the diffusion time embedding and is incorporated through AdaLN.
+In practice, $r_{mlp}$ is set to 4.
+For MMAR-0.5B, $N_{diff}=8$ and $d_{mlp}=1024$, while for MMAR-7B, $N_{diff}=12$ and $d_{mlp}=2048$.
+Image Tokenizer.
+MMAR employs the publicly available KL-16 image tokenizer from LDM [47].
+This tokenizer processes an 256 $\times$ 256 image into 256 image tokens, each with $d=16$ channels, and it is kept frozen during training.
+EmbeddingViT.
+The EmbeddingViT module is implemented with a ViT[11] encoder with $N_{ViT}=16$ layers and $d_{ViT}=1024$ hidden state channels, processing image tokens into visual embeddings with stronger context awareness.
+We integrate a learnable position embedding for each image token in EmbeddingViT, corresponding to its 2D position on the image.
+LLM.
+We initialize our LLM using parameters from the open-source QWen2 series models [63].
+To preserve text modeling capabilities, we keep the LLM parameters fixed during training and add PLoRA [10] as the image expert, where only the image tokens in the input pass through the introduced LoRA [22] adapters.
+The PLoRA is applied for each linear layer within the original LLM.
+Considering the time cost, our ablation study employs QWen2-0.5B-Instruct, with $r=512$ and $\alpha=128$ for PLoRA.
+Furthermore, we use QWen2-7B-Instruct to explore our method’s scale-up capability with $r=1280$ and $\alpha=128$ for PLoRA.
+The corresponding MMAR models are denoted by MMAR-0.5B and MMAR-7B, respectively.
+Figure 6: Details of Diffusion MLP architecture.
+Projector.
+We use two MLP-based projectors [68, 69, 67] to connect different modules’ input and output.
+Specifically, the PostProjector connects the Image Tokenizer output to the EmbeddingViT’s input with a 2-layer MLP, whereas the VisProjector connects EmbeddingViT output to the LLM’s input with a 3-layer MLP.
+Training Task Proportion.
+MMAR models $P(I|T)$, $P(T|I)$ and $P(I)$ through the conditional image generation task, the image understanding task and the unconditional image generation task, respectively.
+Through the first and the second stage training, we set the proportion of these three tasks as 50%, 45%, and 5%, respectively.
+During the optional third training stage, only image generation tasks are applied, and the proportion of the conditional and unconditional image generation tasks are set to 90% and 10%, respectively.
+Mixed-Precision Inference.
+To enhance numerical stability during image generation, particularly when using small step sizes in the DDPM sampling process, we perform the model’s forward pass in bfloat16 (matching the training precision) but cast the output to float32 before DDPM sampling.
+This mitigates potential rounding error without significant computational overhead, improving sampling accuracy efficiently.
+Ablation-VQ.
+Based on the MMAR-0.5B framework, we replace the Image Tokenizer from LDM-KL-16 to LDM-VQ-16.
+The image codes extracted using LDM-VQ-16 are then passed through a projector to increase the channel size to match the LLM’s hidden size.
 Subsequently, we add a decoding Linear layer, which takes the hidden states of the LLM’s output image portion as input and maps them to the image codebook.
-The Cross Entropy loss is then calculated between these mapped values and the actual VQ codes.
-Transfusion
-Following the theoretical ideas presented in the Transfusion paper, we adopt a simple linear mapping.
-After extracting image tokens using KL-16, if the task is image first, we add noise within a 500-time step to the image tokens.
-Otherwise, we add noise within a 1000-step.
-After the linear mapping, we add the image token a learnable time embedding corresponding to the time step as input to the LLM.
-We also maintain the bidirectional attention mechanism.
-After passing through the LLM, we first map the output back to the original token channel count using a linear layer and compute the MSE loss for the predicted noise.
-During generation, we treat LLM as a denoised model, with the condition being the concatenation of the text and the image tokens to be generated.
-We adopt the learning approach of Transfusion but conduct experiments based on our training tasks and stage divisions.
-Projector
-In order to accomplish channel alignment, we introduced two Projectors designed to scale the channels.
-Both Projectors consist of a simple linear layer and multiple blocks composed of activation layers and linear layers.
-The PostProjector comprises one block, whereas the VisProjector contains two blocks.
-A.2 Minimizing the Numerical Error in Diffusion Models
+The Cross Entropy loss is then calculated between these mapped values and the ground-truth VQ codes.
+Transfusion Reimplementation.
+Our re-implementation version of Transfusion shares most of the MMAR-0.5B’s model architecture except for the processing of the model input and output.
+Following the Transfusion paper, we adopt a input linear projection to convert the output of LDM-KL-16 image tokenizer into the LLM input representation, and use an output linear projection to convert the LLM output into the predicted noise.
+Noise of different levels are added to the image token input according to different pretraining tasks.
+For the image-to-text task, the diffusion time step is uniformly sampled within $t\in[0,500]$, while for the text-to-image task, the diffusion time step is uniformly sampled within $t\in[0,1000]$.
+A learnable time embedding corresponding to the time step $t$ is added after the input linear projection.
+MSE loss is calculated between the predicted and ground-truth noise.
+During inference, we treat the LLM as a diffusion model, with the condition being the concatenation of the text and the noisy image tokens of the previous diffusion time step $t+1$.
+6.2 Minimizing the Numerical Error in Diffusion Models
 To make our discussion clearer, we switch the diffusion noise schedule into an angular form as follows:
@@
 $\begin{cases}x=\frac{\sin\psi_{t}x^{(t)}-\sin\phi_{t}u^{(t)}/r_{t}}{\cos\phi_{t}\sin\psi_{t}-\cos\psi_{t}\sin\phi_{t}}=\frac{\sin\psi_{t}x^{(t)}-\sin\phi_{t}u^{(t)}/r_{t}}{\sin(\psi_{t}-\phi_{t})},\\ \epsilon=\frac{\cos\psi_{t}x^{(t)}-\cos\phi_{t}u^{(t)}/r_{t}}{\sin\phi_{t}\cos\psi_{t}-\sin\psi_{t}\cos\phi_{t}}=-\frac{\cos\psi_{t}x^{(t)}-\cos\phi_{t}u^{(t)}/r_{t}}{\sin(\psi_{t}-\phi_{t})}.\end{cases}$ (12)
-Now we consider the general form of DDIM sampling step (Song et al., 2021):
+Now we consider the general form of DDIM sampling step [51]:
 $x^{(t-1)}=\cos\phi_{t-1}\hat{x}_{\theta}(x^{(t)})+\sin\phi_{t-1}\hat{\epsilon}_{\theta}(x^{(t)}),$ (13)
@@
 Therefore, we can further represent $x^{(t-1)}$ in the following form:
-$x^{(t-1)}$ $=\cos\phi_{t-1}\frac{\sin\psi_{t}x^{(t)}-\sin\phi_{t}u_{\theta}^{(t)}/r_{t}}{\sin(\psi_{t}-\phi_{t})}-\sin\phi_{t-1}\frac{\cos\psi_{t}x^{(t)}-\cos\phi_{t}u_{\theta}^{(t)}/r_{t}}{\sin(\psi_{t}-\phi_{t})}$
+$x^{(t-1)}=$ $\cos\phi_{t-1}\frac{\sin\psi_{t}x^{(t)}-\sin\phi_{t}u_{\theta}^{(t)}/r_{t}}{\sin(\psi_{t}-\phi_{t})}$
+$-\sin\phi_{t-1}\frac{\cos\psi_{t}x^{(t)}-\cos\phi_{t}u_{\theta}^{(t)}/r_{t}}{\sin(\psi_{t}-\phi_{t})}$
 $=$ $\frac{\sin(\phi_{t-1}-\phi_{t})u_{\theta}^{(t)}/r_{t}-\sin(\phi_{t-1}-\psi_{t})x^{(t)}}{\sin(\psi_{t}-\phi_{t})}.$ (14)
@@
 As a result, the total step number of DDIM sampling is increased proportionally, which cancels out the effect of error reduction of each single step.
-The factor $u_{\theta}^{(t)}/r_{t}$ is not adjustable since it constantly has a unit standard deviation.
-This can be verified by the following calculation:
-$\mathbb{E}[(u^{(t)}/r_{t})^{2}]=\mathbb{E}[(\cos\psi_{t}x+\sin\psi_{t}\epsilon)^{2}]=\cos^{2}\psi_{t}\mathbb{E}[x^{2}]+\sin^{2}\psi_{t}.$ (19)
-In common practice, image tokens $x$ are normalized into unit standard deviation. Therefore, $\mathbb{E}[(u^{(t)}/r_{t})^{2}]=\cos^{2}\psi_{t}+\sin^{2}\psi_{t}=1$.
+The factor $u_{\theta}^{(t)}/r_{t}$, which represents the normalized model prediction, is largely decided by the optimization progress and the target data distribution.
+An approximation can be derived from a perfect prediction, that is $u_{\theta}^{(t)}/r_{t}=u^{(t)}/r_{t}$.
+In this way,
+$\mathbb{E}[(u_{\theta}^{(t)}/r_{t})^{2}]$ $\approx\mathbb{E}[(\cos\psi_{t}x+\sin\psi_{t}\epsilon)^{2}]$
+$=\cos^{2}\psi_{t}\mathbb{E}[x^{2}]+\sin^{2}\psi_{t}.$ (19)
+In common practice, image tokens $x$ are normalized into unit standard deviation, leading to $\mathbb{E}[(u_{\theta}^{(t)}/r_{t})^{2}]\approx\cos^{2}\psi_{t}+\sin^{2}\psi_{t}=1$.
+This is a constant number, meaning that there is not much potential to reduce the total numerical error via tuning down $u_{\theta}^{(t)}/r_{t}$.
 If we decide to scale up our model, it is better to leverage the pre-trained LLMs as well as the highly efficient training infrastructure that is specifically optimized for LLMs.
@@
 A simple solution is to set $\psi_{t}-\phi_{t}=\pi/2$, resulting in the following parameterization:
-$u^{(t)}=r_{t}\cos(\phi_{t}+\pi/2)x+r_{t}\sin(\phi_{t}+\pi/2)\epsilon=r_{t}(\cos\phi_{t}\epsilon-\sin\phi_{t}x).$ (20)
+$u^{(t)}$ $=r_{t}\cos(\phi_{t}+\pi/2)x+r_{t}\sin(\phi_{t}+\pi/2)\epsilon$
+$=r_{t}(\cos\phi_{t}\epsilon-\sin\phi_{t}x).$ (20)
 Note that $r_{t}$ is still undetermined, which reflects the scale of $u^{(t)}$.
@@
 $u^{(t)}=\cos\phi_{t}\epsilon-\sin\phi_{t}x.$ (21)
-We notice that this parameterization is coincidentally the “v-prediction” parameterization (Salimans & Ho, 2022).
+We notice that this parameterization is coincidentally the “v-prediction” parameterization [49].
 Note that, however, “v-prediction” is initially proposed for the efficient distillation of diffusion models, rather than reducing the numerical error of diffusion models.
 To the best of our knowledge, our work is the first to derive “v-prediction” parameterization from the first principle of minimizing the numerical error in diffusion models.
-A.3 Deriving Theoretical Numerical Error for $\epsilon$-Prediction Models
+Table 6: Detailed visual understanding evaluation results.
+MMAR-0.5B MMAR-0.5B MMAR-7B
+Benchmark Chameleon-7B Transfusion* Show-o MMAR-0.5B w/ $\epsilon$-pred. w/ VQ MMAR-7B w/ Stg3
+AI2D [24] 34.81 40.22 32.48 43.43 41.90 41.90 64.64 63.54
+ChartQA [40] 3.84 9.56 11.32 10.20 10.36 9.36 13.64 12.52
+DocVQA [41] 1.51 6.72 18.24 7.62 6.77 6.79 11.12 10.62
+Hallu.Bench [18] 39.01 41.54 40.90 42.80 41.11 41.54 53.10 52.05
+MathVista [39] 21.90 22.60 23.20 21.60 23.10 22.90 32.40 31.30
+MMBenchCN [35] 10.14 27.23 0.52 43.99 38.83 31.87 70.53 66.75
+MMBenchEN [35] 13.32 29.47 42.44 48.45 45.53 37.54 70.45 67.78
+MMEP [15] 125.8 594.3 1182.7 882.1 880.7 618.2 1486.9 1421.9
+MMEC [15] 33.9 206.1 225.0 256.8 232.1 273.2 268.9 303.6
+MMMU [66] 24.00 29.33 26.44 29.33 25.33 29.67 41.33 47.33
+MMStar [5] 20.47 28.13 32.00 32.13 31.07 28.07 41.87 40.87
+MMVet [65] 7.34 13.90 20.87 18.49 17.98 14.45 30.64 29.17
+OCRBench [36] 0.50 2.30 15.20 18.70 7.10 2.10 25.80 21.20
+POPE [31] 30.86 66.90 84.50 70.74 71.14 66.98 84.02 83.21
+RealWorldQA 27.06 36.99 27.97 38.30 35.16 36.60 53.59 52.68
+ScienceQA [38] 44.83 45.92 41.82 47.54 45.21 45.35 74.39 73.20
+SEEDBench [26] 34.61 42.40 51.61 55.70 53.72 44.93 68.63 66.59
+TextVQA [50] 5.43 9.94 38.35 16.77 12.40 9.46 24.37 21.35
+Average 18.34 28.26 33.06 34.56 32.21 29.70 48.25 47.18
+6.3 Deriving Theoretical Numerical Error for $\epsilon$-Prediction Models
 The $\epsilon$-prediction parameterization corresponds to $\psi_{t}=\frac{\pi}{2}$ in the angular parameterization form given by Eq.11.
@@
 In order to calculate the equivalent $v^{(t)}_{\theta}$ value, we need to represent $v^{(t)}_{\theta}$ with the predicted $\epsilon_{\theta}$ and the known $x^{(t)}$, which is calculated as follows:
-$v^{(t)}_{\theta}=\cos\phi_{t}\epsilon_{\theta}-\sin\phi_{t}\hat{x}_{\theta}(x^{(t)})=\cos\phi_{t}\epsilon_{\theta}-\sin\phi_{t}\frac{x^{(t)}-\sin\phi_{t}\epsilon_{\theta}}{\cos\phi_{t}}=\frac{\epsilon_{\theta}}{\cos\phi_{t}}-\tan\phi_{t}x^{(t)}.$ (24)
+$v^{(t)}_{\theta}$ $=\cos\phi_{t}\epsilon_{\theta}-\sin\phi_{t}\hat{x}_{\theta}(x^{(t)})$
+$=\cos\phi_{t}\epsilon_{\theta}-\sin\phi_{t}\frac{x^{(t)}-\sin\phi_{t}\epsilon_{\theta}}{\cos\phi_{t}}$
+$=\frac{\epsilon_{\theta}}{\cos\phi_{t}}-\tan\phi_{t}x^{(t)}.$ (24)
 Considering the numerical error, we get:
@@
 Taking numerical error effect into the v-prediction-based diffusion loss, we get:
-$\mathbb{E}[(v^{(t)}-\tilde{v}^{(t)}_{\theta})^{2}]=\mathbb{E}[(v^{(t)}-v^{(t)}_{\theta}-e^{(t)})^{2}]=\mathbb{E}[(v^{(t)}-v^{(t)}_{\theta})^{2}]-2\mathbb{E}[(v^{(t)}-v^{(t)}_{\theta})e^{(t)}]+\mathbb{E}[(e^{(t)})^{2}].$ (26)
+$\mathbb{E}[(v^{(t)}-\tilde{v}^{(t)}_{\theta})^{2}]=$ $\mathbb{E}[(v^{(t)}-v^{(t)}_{\theta}-e^{(t)})^{2}]$
+$=$ $\mathbb{E}[(v^{(t)}-v^{(t)}_{\theta})^{2}]-2\mathbb{E}[(v^{(t)}-v^{(t)}_{\theta})e^{(t)}]$
+$+\mathbb{E}[(e^{(t)})^{2}].$ (26)
 Due to the fact that numerical error $e^{(t)}$ is independent from the training loss and that the expectation of $e^{(t)}$ is 0, we get $\mathbb{E}[(v^{(t)}-v^{(t)}_{\theta})e^{(t)}]=0$.
@@
 This is the theoretical numerical error of the v-prediction diffusion loss for an $\epsilon$-prediction model.
-A.4 CFG With $v$-prediction
+Figure 8: The impact of CFG scale on image generation quality.
+6.4 CFG With $v$-prediction
 From Equation $v^{(t)}_{i}=\sqrt{\bar{\alpha}_{t}}\mathbf{\epsilon}-\sqrt{1-\bar{\alpha}_{t}}x_{i}$, we can derive the following equation.
@@
 Ultimately, we obtain $v=v_{u}+\omega(v_{c}-v_{u})$. The CFG of $v$ and $\epsilon$ are equivalent.
-A.5 Examples:The effect of the stage 2 training
-Figure 5: The impact of the second training stage on image generation capability.
-A.6 Examples:Image Generation
-Figure 6: Generated images from MMAR-7B
-A.7 Detailed Visual Understanding Evaluation Results
-A total of 18 visual understanding benchmarks from VLMEvalKit (Duan et al., 2024) are used to evaluate MMAR models comprehensively.
+Figure 9: Generated images from MMAR-7B after the third training stage (1/2).
+6.5 Impact of CFG Scaling
+We select models from the second and fourth epochs of the first stage as starting points for the second stage, train them for 3 epochs, and then test the MSCOCO FID-30K under varying CFG intensities.
+As shown in Fig.
+8, our method achieves better FID scores as the CFG scale increases from 1 to 10.
+It is worth noting that most probabilistic generative models typically have a CFG scale between 1.5 and 5.
+Additionally, it is observed that a longer training duration in the first stage consistently results in better generation outcomes at all CFG scales.
+6.6 Detailed Visual Understanding Evaluation Results
+A total of 18 visual understanding benchmarks from VLMEvalKit [12] are used to evaluate MMAR models comprehensively.
 The evaluation is also conducted on the existing joint image-text probabilistic models using the publicly available checkpoints [^3] https://huggingface.co/facebook/chameleon-7b [^4] https://huggingface.co/showlab/show-o-w-clip-vit.
@@
 The average score is calculated on the normalized score of all the benchmarks including MME.
-Table 4: Detailed visual understanding evaluation results.
-Benchmark Chameleon-7B Transfusion* Show-o MMAR-0.5B MMAR-0.5B w/ $\epsilon$-pred. MMAR-0.5B w/ VQ MMAR-7B
-AI2D 34.81 40.22 32.48 43.43 41.90 41.90 62.63
-ChartQA 3.84 9.56 11.32 10.20 10.36 9.36 12.72
-DocVQA 1.51 6.72 18.24 7.62 6.77 6.79 10.28
-Hallu.Bench 39.01 41.54 40.90 42.80 41.11 41.54 50.47
-MathVista 21.90 22.60 23.20 21.60 23.10 22.90 32.10
-MMBenchCN 10.14 27.23 0.52 43.99 38.83 31.87 66.49
-MMBenchEN 13.32 29.47 42.44 48.45 45.53 37.54 66.32
-MMEP 125.8 594.3 1182.7 882.1 880.7 618.2 1393.9
-MMEC 33.9 206.1 225.0 256.8 232.1 273.2 320.0
-MMMU 24.00 29.33 26.44 29.33 25.33 29.67 42.78
-MMStar 20.47 28.13 32.00 32.13 31.07 28.07 40.00
-MMVet 7.34 13.90 20.87 18.49 17.98 14.45 27.80
-OCRBench 0.50 2.30 15.20 18.70 7.10 2.10 23.10
-POPE 30.86 66.90 84.50 70.74 71.14 66.98 83.02
-RealWorldQA 27.06 36.99 27.97 38.30 35.16 36.60 52.42
-ScienceQA 44.83 45.92 41.82 47.54 45.21 45.35 71.05
-SEEDBench 34.61 42.40 51.61 55.70 53.72 44.93 64.52
-TextVQA 5.43 9.94 38.35 16.77 12.40 9.46 22.05
-Average 18.34 28.26 33.06 34.56 32.21 29.70 46.52
-A.8 Analysis on the Visual Understanding Ability of Transfusion-Like Methods
-In this section, we will quantitatively analyze the incomplete information usage problem in transfusion-like methods.
-The experiments are conducted using our re-implemented transfusion model.
-The model is evaluated on MMBench using different noise-level image input.
-Specifically, we take the diffusion time-step of 0, 125, 250, 375, 500, and 750 for testing.
-Among the MMBench tasks, we focus on tasks whose performance is significantly affected by noise-levels.
-Fig.
-7 depict the performance trend of these tasks across different noise levels.
-Figure 7: The trend of MMBench tasks’ performance with respect to input image noise levels.
-(A): The best performance of most tasks are achieved at a non-zero noise level.
-(B): The object localization performance consistently gets higher as the noise level grows up.
-Intuitively, increasing noise levels should lead to a consistent performance decline across all visual understanding tasks due to information loss.
-However, our findings reveal a more nuanced relationship.
-We find that the best performance of most visual understanding tasks are achieved at a non-zero noise level.
-Especially, the object localization task, which is more focused on the image layout information, consistently performs better as the noise level grows higher.
-This suggests that transfusion-like methods, when limited to a single noise level input, fail to fully exploit the information of the input image.
-Instead, visual information appears to be distributed across different noise levels within the model, with the input noise level dictating which information is ultimately utilized.
+6.7 Examples:Image Generation Sample
+In Fig.910, we showcase some generated examples from MMAR-7B after the third stage training, featuring animals, plants, real-world scenes, artistic scenes, and counterfactual themes.
+Figure 10: Generated images from MMAR-7B after the third training stage (2/2).
+6.8 Numerical Error Empirical Validation
+Figure 11: Diffusion sampling results with different prediction targets and training precision. Figures are scaled differently to include all generated samples.
+We provide direct empirical demonstration using a 2D chessboard distribution (class-conditioned quadrant mapping).
+As Fig.11 reveals: (1) v-prediction maintains stability even with bf16 precision at scale (1M samples, Fig.(a)), achieving $<10^{-6}$ token error rate.
+(2) $\epsilon$-prediction exhibits visible artifacts (Fig.(b),(c)) with 0.2% token error rate (bf16), equivalent to 51.2% image-level failure for 256-token images.
+(3) Precision elevation (fp32) reduces artifacts to 0.04%.
+These observations validate v-prediction’s critical role in mitigating low-precision training risks.